### v2

# EPW varieties as moduli spaces on ordinary GM surfaces and special GM threefolds

Abstract.
We show that the double dual EPW sextic and double EPW sextic associated with a strongly smooth Gushel–Mukai surface can be realized as moduli spaces of semistable objects with respect to a stability condition on the bounded derived category of it. Also, we observe that double dual EPW surface and double EPW surface associated with a special Gushel–Mukai threefold can be realized as moduli spaces of semistable objects on its Kuznetsov component. As an application, we refine a statement of Bayer and Perry about Gushel–Mukai threefolds with equivalent Kuznetsov components for the special ones.
Key words and phrases:
Guehsl–Mukai varieties, moduli space, derived category
2020 Mathematics Subject Classification:
Primary 18G80 14J28 14J45
The first author is a member of the INdAM group GNSAGA (2025). The second author is supported by SUSY Starting Grant No. 34000-12256019.
1. Introduction
Moduli spaces of semistable sheaves provide new methods to study the geometry of low dimension varieties and provide models to some higher dimensional varieties. Here we refer the readers to Huybrechts and Lehn [26] for a comprehensive account of this theory. On the other hand, moduli spaces of semistable sheaves on special members in a family of varieties are barely understood even for families of surfaces. Nevertheless, moduli spaces of semistable sheaves are in general only compatible with étale covers but not branched ones. These two disadvantages can be partly solved by considering instead the moduli spaces of semistable objects with respect to a notion of stability conditions introduced by Bridgeland [10].
In this article, we will study two moduli spaces associated with a large family of K3 surfaces called the strongly smooth ordinary Gushel–Mukai surfaces and two moduli spaces on the natural double covers branched over these K3 surfaces called the special Gushel–Mukai threefolds.
Throughout the article, everything will be over the field of complex numbers, categories are essentially small, and functors are additive.
1.1. GM varieties
An ordinary Gushel–Mukai surface is a smooth dimension transverse intersection $S=\operatorname{Gr}(2,V_{5})\cap\mathbb{P}(W)\cap Q$ where $V_{5}$ is a five-dimensional vector space, $W\subset\mathsf{\Lambda}^{2}V_{5}$ is a seven-dimensional vector subspace, and $Q\subset\mathbb{P}(W)$ is a quadratic hypersurface, such that the intersection $M_{S}:=\operatorname{Gr}(2,V_{5})\cap\mathbb{P}(W)$ is transversal. The surface $S$ is a Brill–Noether general K3 surface of degree $10$ under the ample polarization $H$ given by $\mathscr{O}_{S}(H)\mathscr{O}_{\mathbb{P}(W)}(1)|_{S}$.
The surface $S$ is called strongly smooth if the intersection $M_{S}$ is smooth. The double cover $X\rightarrow M_{S}$ branched over a strongly smooth ordinary Gushel–Mukai surface $S$ is called the special Gushel–Mukai threefold associated with $S$. Such a threefold $X$ has Picard rank $1$ and is Fano of index one and degree $10$ under the polarization $\mathscr{O}_{X}(1)$ given by pulling back $\mathscr{O}_{\mathbb{P}(W)}(1)$.
The covering involution on $X$ induces a $\mathfrak{S}_{2}$-group action on the Kuznetsov component
${\textbf{{Ku}}}(X)=\{E\in{\textbf{{D}}}^{b}(X)\,|\,\operatorname{Hom}(\mathscr{U}_{X}^{\vee},E[n])=\operatorname{Hom}(\mathscr{O}_{X},E[n])=0\textup{ for any }n\in\mathbb{Z}\}$
of $X$, where the vector bundle $\mathscr{U}_{X}$ is induced by the tautological bundle $\mathscr{U}_{G}$ on $\operatorname{Gr}(2,V_{5})$. The associated equivariant category is equivalent to ${\textbf{{D}}}^{b}(S)$ according to [30, Section 8.2]. The dual cyclic group $\mathfrak{S}^{\vee}_{2}$ acts on ${\textbf{{D}}}^{b}(S)$ by an involutive autoequivalence $\Pi$ on ${\textbf{{D}}}^{b}(X)$ and the associated equivariant category is reversely equivalent to ${\textbf{{Ku}}}(X)$ by [19, Theorem 1.3].
According to [33, Secion 6.1], one can find a stability condition $\sigma_{S}$ on ${\textbf{{D}}}^{b}(S)$ for any strongly smooth ordinary Gushel–Mukai surface $S$ such that the moduli spaces of semistable objects with respect to $\sigma_{S}$ detect many geometry of $S$. Moreover, the stability condition $\sigma_{S}$ is compatible with the $\mathfrak{S}^{\vee}_{2}$-group action on ${\textbf{{D}}}^{b}(S)$ so that it induces a stability condition $\sigma_{X}$ on the equivariant category ${\textbf{{Ku}}}(X)$ and vice versa by virtue of [43, Section 4.3].
The forget functors between ${\textbf{{D}}}^{b}(S)$ and ${\textbf{{Ku}}}(X)$ induce covering morphisms between moduli spaces of semistable objects with respect to $\sigma_{S}$ and $\sigma_{X}$. The properties and first examples of these covering morphisms have been investigated in [33, 43]. In particular, some classic branched double covers can be recovered from these examples. This article is aim to recover some EPW varieties associated with $S$ and $X$ as moduli spaces, expanding results in [33, Section 6].
1.2. EPW varieties
The vector space $V_{6}:=H^{0}(\mathbb{P}(W),\mathscr{I}_{S}(2))$ of quadrics in $W$ containing an ordinary Gushel–Mukai surface $S$ is six-dimensional and each ordinary Gushel–Mukai surface $S$ is canonically associated with a Lagrangian subspace $A(S)\subset\mathsf{\Lambda}^{3}V_{6}$. One can construct a four-dimensional projective variety $\tilde{Y}^{\geq 1}_{A(S)}$ from $A(S)$, called the double EPW sextics associated with the Gushel–Mukai surface $S$. The annihilator $A(S)^{\perp}\subset\mathsf{\Lambda}^{3}V^{\vee}_{6}$ of the Lagrangian subspace $A(S)$ is also a Lagrangian subspace and gives another four-dimension projective variety $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}$ called the double dual EPW sextics associated with $S$.
The singular locus of $\tilde{Y}^{\geq 1}_{A(S)}$ is an integral normal Cohen–Macaulay surface $Y^{\geq 2}_{A(S)}$ called the EPW surface associated with $S$. The unique double cover over $Y^{\geq 2}_{A(S)}$ branched on its singular locus is called the double EPW surface and denoted by $\tilde{Y}^{\geq 2}_{A(S)}$. Similarly, one has the associated dual EPW surface $Y^{\geq 2}_{A(S)^{\perp}}$ and associated double dual EPW surface $\tilde{Y}^{\geq 2}_{A(S)^{\perp}}$.
The special Gushel–Mukai threefold $X$ over a strongly smooth ordinary Gushel–Mukai surface $S$ also induces a Lagrangian subspace $A(X)\subset\mathsf{\Lambda}^{3}V_{6}$ which is canonically isomorphic to the subspace $A(S)$. So the notations $A(X)$ and $A(S)$ will be used interchangeably.
1.3. Main results
In [33], the sextics $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}$ and $\tilde{Y}^{\geq 1}_{A(S)}$ are realized as moduli spaces of semistable objects with respect to $\sigma_{S}$ for a general strongly smooth ordinary Gushel–Mukai surface $S$ based on the descriptions in [37, 40]. Here we provide the modular description of $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}$ for any strongly smooth ordinary Gushel–Mukai surface following [18, Section 7.5].
Theorem 1.1.
The moduli space $\textsf{{M}}_{\sigma_{S}}(v_{1})$ of semistable objects with Mukai vector $v_{1}=(1,0,-1)$ is isomorphic to the double dual EPW sextics $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}$ associated $S$.
However, a useful birational model for $\tilde{Y}^{\geq 1}_{A(S)}$ is not available. In this case, one can pass to a strongly smooth ordinary Gushel–Mukai surface $S^{\prime}$ which is derived equivalent to the surface $S$ under a $\mathfrak{S}_{2}^{\vee}$-equivariant equivalence and satisfies $\tilde{Y}_{A(S)}\cong\tilde{Y}_{A(S^{\prime})^{\perp}}$ using [7, 40]. Then one can deduce the following statement using the results in [33, Section 6].
Proposition 1.2.
The moduli space $\textsf{{M}}_{\sigma_{S}}(v_{2})$ of semistable objects with Mukai vector $v_{2}=(2,-H,2)$ is isomorphic to the double EPW sextics $\tilde{Y}^{\geq 1}_{A(S)}$ associated $S$.
Consequently, the inflation functor ${\textbf{{Ku}}}(X)\rightarrow{\textbf{{D}}}^{b}(S)$ induces two double covers
$\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})\rightarrow Y^{\geq 2}_{A(S)^{\perp}}\quad\textup{and}\quad\textsf{{M}}_{\sigma_{X}}(-\kappa_{2})\rightarrow Y^{\geq 2}_{A(S)}$
branched over the singular loci of the targets, where $\kappa_{1}$ and $\kappa_{2}$ are classes in the numerical Grothendieck group $\textsf{{N}}({\textbf{{Ku}}}(X))$ such that $\textup{{ch}}(\kappa_{1})=1-h^{2}/5$ and $\textup{{ch}}(\kappa_{2})=2-h+h^{3}/12$ for the class $h:=\textup{{ch}}_{1}(\mathscr{O}_{X}(1))$. One can see $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})\cong\tilde{Y}^{\geq 2}_{A(S)^{\perp}}$ from the description for the moduli space in [28] and for the EPW variety in [18], and the following statement by [33].
Proposition 1.3.
The double cover $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})\rightarrow Y^{\geq 2}_{A(S)^{\perp}}$ coincides with $\tilde{Y}^{\geq 2}_{A(S)^{\perp}}\rightarrow Y^{\geq 2}_{A(S)^{\perp}}$.
Similar to our identification for $\tilde{Y}^{\geq 1}_{A(S)}$, one can show the following statement.
Proposition 1.4.
One has $\textsf{{M}}_{\sigma_{X}}(-\kappa_{2})\cong\tilde{Y}^{\geq 2}_{A(X)}$ for any special Gushel–Mukai threefold.
The classes $\kappa_{1}$ and $\kappa_{2}$ form a basis for $\textsf{{N}}({\textbf{{Ku}}}(X))$ which is a rank two free abelian group so that an equivalence ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$ either preserves or exchanges them up to a sign. It is conjectured in [31] and proved in [7, Theorem 5.12] using homological projective duality that two special Gushel–Mukai threefolds $X$ and $X^{\prime}$ have equivalent Kuznetsov components if and only if one has $A(X)\cong A(X^{\prime})$ or $A(X)\cong A(X^{\prime})^{\perp}$. Using the computations in [7, 33], a refinement can be established as an application of Theorem 1.1 and Proposition 1.2.
Corollary 1.5.
Consider two special Gushel–Mukai threefolds $X$ and $X^{\prime}$, then $A(X)\cong A(X^{\prime})$ if and only if one has ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$ sending $\kappa_{1}$ to $\kappa^{\prime}_{1}$ and $\kappa_{2}$ to $\kappa^{\prime}_{2}$ up to a sign; $A(X)^{\perp}\cong A(X^{\prime})$ if and only if one has ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$ sending $\kappa_{1}$ to $\kappa^{\prime}_{2}$ and $\kappa_{2}$ to $\kappa^{\prime}_{1}$ up to a sign.
Remark 1.6.
Suppose in addition that $S$ does not contain any conic or lines, then ${\textbf{{Ku}}}(X)$ determines the special Gushel–Mukai threefold $X$. This special case has been proved in [28].
In the end of this article, we exploit our results together with the Brill–Noether locus identification of a special Gushel–Mukai threefold in [29] to investigate the special Gushel–Mukai threefolds sharing the same Lagrangian subspaces.
2. GM Varieties and EPW Varieties
In this section, some properties of Gushel–Mukai varieties and the EPW varieties will be recalled. The interested readers can find more information about Gushel–Mukai varieties in the series of articles [13, 14, 15, 16, 18] and EPW varieties in [17, 37, 38, 39, 40, 41].
2.1. The Gushel–Mukai varieties
In the introduction, we have defined the strongly smooth ordinary Gushel–Mukai surfaces and the special Gushel–Mukai threefolds. Here we will provide the general definition. Let $V_{5}$ be a five-dimensional vector space, and consider the Plücker embedding $\operatorname{Gr}(2,V_{5})\subset\mathbb{P}(\mathsf{\Lambda}^{2}V_{5})$. Let $K$ be an arbitrary vector space, consider the cone
$\operatorname{CGr}(2,V_{5})\subset\mathbb{P}(\mathsf{\Lambda}^{2}V_{5}\oplus K)$
with vertex $\mathbb{P}(K)$. Choose a subspace $W\subset\mathsf{\Lambda}^{2}V_{5}\oplus K$ and a quadratic hypersurface $Q\subset\mathbb{P}(W)$.
Definition 2.1.
A Gushel–Mukai variety $X$ is a smooth $n$-dimensional intersection
$X=\operatorname{CGr}(2,V_{5})\cap\mathbb{P}(W)\cap Q$
for $2\leq n\leq 6$, such that the dimension of the vector space $W$ is $n+5$.
A Gushel–Mukai variety $X$ is canonically polarized by restriction of the hyperplane class on $\mathbb{P}(W)$. The ample polarization $H$ satisfies $H^{n}=10$ and $K_{X}=-(n-2)H$.
Example 2.2.
A polarized Gushel–Mukai surface $(S,H)$ is a Brill–Noether general K3 surface of degree $10$. It means that $H$ is very ample, and there are no cubic elliptic curves on $S$.
Example 2.3.
The Picard group of a Gushel–Mukai threefold $X$ is generated by the canonical ample polarization $H$, and $H$ makes $X$ a Fano threefold of index $1$ and degree $10$.
The intersection $M_{X}=\operatorname{CGr}(2,V_{5})\cap\mathbb{P}(W)$ is called the Grassmannian hull of the Gushel–Mukai variety $X$. A Gushel–Mukai variety is called strongly smooth if its Grassmannian hull is smooth. Any Gushel–Mukai variety with dimension $n\geq 3$ is strongly smooth and the strongly smooth Gushel–Mukai surfaces are characterized by the next statement.
Proposition 2.4 ([20, Lemma 2.7, Lemma 2.8]).
A Gushel–Mukai surface $(S,H)$ is strongly smooth if and only if it does not have any divisor $E$ satisfying $E^{2}=0$ and $E\cdot H=4$.
In fact, due to [8, Remark 2.19], a Gushel–Mukai surface is possibly strongly smooth only when it is ordinary in the following definition.
Definition 2.5.
A Gushel–Mukai variety $X$ is called ordinary if the vertex of $\operatorname{CGr}(2,V_{5})$ is not contained in $M_{X}$ and is called special if the vertex of $\operatorname{CGr}(2,V_{5})$ is contained in $M_{X}$.
An $(n+1)$-dimensional special Gushel–Mukai variety $X$ admits a branched double cover over the Grassmannian hull $M_{Y}$ of a strongly smooth ordinary Gushel–Mukai variety $Y$ with dimension $n$ such that the branched locus is $Y\subset M_{Y}$. Conversely, the double cover over any Grassmannian hull $M_{Y}$ branched on the $n$-dimensional strongly smooth ordinary Gushel–Mukai variety $Y\subset M_{Y}$ is a special Gushel–Mukai variety with dimension $n+1$.
In this case, the special Gushel–Mukai variety $X$ is called the opposite Gushel–Mukai variety of the ordinary Gushel–Mukai variety $Y$ and vice versa.
Remark 2.6.
Even though a special Gushel–Mukai variety $X$ does not admit an embedding into $\operatorname{Gr}(2,V_{5})$ as for the ordinary ones, one still has a natural morphism $X\rightarrow\operatorname{Gr}(2,V_{5})$ which induces the double cover branched over the opposite ordinary Gushel–Mukai surface.
2.2. The Lagrangian subspaces and EPW varieties
Any Gushel–Mukai variety
$X=\operatorname{CGr}(2,V_{5})\cap\mathbb{P}(W)\cap Q$
is canonically associated with a Lagrangian subspace $A(X)\subset\mathsf{\Lambda}^{3}V_{6}$ for $V_{6}=H^{0}(\mathbb{P}(W),\mathscr{I}_{X}(2))$ according to [15, Theorem 3.10]. One can define a series of projective varieties associated with the Lagrangian subspace. These varieties are introduced and intensively studied by O’Grady.
In general, let $V_{6}$ be a $6$-dimensional vector space and $A\subset\mathsf{\Lambda}^{3}V_{6}$ be the Lagrangian subspace for the $\det(V_{6})$-valued symplectic form $\omega$ defined by $\omega(\zeta,\eta)=\zeta\land\eta$, then the subset
$\{[v]\in\mathbb{P}(V_{6})\,|\,\dim(A\cap\left(v\land\mathsf{\Lambda}^{2}V_{6}\right))\geq\ell\}$
can be endowed with a scheme structure $Y_{A}^{\geq\ell}$ for any integral $\ell$ according to [37, Section 2].
The subscheme $Y_{A}^{\geq 1}\subset\mathbb{P}(V_{6})$ can a priori be the whole space $\mathbb{P}(V_{6})$. To avoid this, one needs an additional condition.
Definition 2.7.
A non-zero element of a Lagrangian subspace $A\subset\mathsf{\Lambda}^{3}V_{6}$ is called decomposable once it can be written as the form $v_{1}\land v_{2}\land v_{3}$.
The Lagrangian subspace associated with a strongly smooth Gushel–Mukai variety does not contain any decomposable element.
Theorem 2.8 (O’Grady).
Consider a Lagrangian subspace $A\subset\mathsf{\Lambda}^{3}V_{6}$ which does not contain decomposable elements, then
  • (a)
    $Y_{A}^{\geq 1}$ is an integral normal sextic hypersurface in $\mathbb{P}(V_{6})$;
  • (b)
    $Y_{A}^{\geq 2}=\operatorname{Sing}(Y_{A}^{\geq 1})$ and is an integral normal Cohen–Macaulay surface of degree $40$;
  • (c)
    $Y_{A}^{\geq 3}=\operatorname{Sing}(Y_{A}^{\geq 2})$ is finite and smooth;
  • (d)
    $Y_{A}^{\geq 4}=\varnothing$.
The sextic hypersurface $Y_{A}^{\geq 1}\subset\mathbb{P}(V_{6})$ is called the EPW sextic associated with $A$.
Theorem 2.9 (O’Grady).
Consider a Lagrangian subspace $A\subset\mathsf{\Lambda}^{3}V_{6}$ which does not contain decomposable elements and the associated EPW sextic $Y^{\geq 1}_{A}\subset\mathbb{P}(V_{6})$, then there exists a double covering $f_{A}\colon\tilde{Y}^{\geq 1}_{A}\rightarrow Y_{A}^{\geq 1}$ branched over $Y_{A}^{\geq 2}$ such that
  • (a)
    the variety $\tilde{Y}_{A}^{\geq 1}$ is simply connected;
  • (b)
    the singular locus of is the finite set $f^{-1}_{A}(Y_{A}^{\geq 3})$;
  • (c)
    the variety is a $K3[2]$-type hyperkähler manifold once $Y_{A}^{\geq 3}=\varnothing$.
The variety $\tilde{Y}_{A}^{\geq 1}$ is called the double EPW sextic associated with the Lagrangian subspace $A$.
One of the most important property is that the double EPW sextic can be determined by its associated Lagrangian subspace.
Theorem 2.10 (O’Grady).
Consider two Lagrangian subspaces $A\subset\mathsf{\Lambda}^{3}V_{6}$ and $A^{\prime}\subset\mathsf{\Lambda}^{3}V^{\prime}_{6}$ such that at least one of them contains no decomposable elements, then any isomorphism $\tilde{Y}^{\geq 1}_{A}\rightarrow\tilde{Y}^{\geq 1}_{A^{\prime}}$ induces a linear isomorphism $\varphi\colon V_{6}\rightarrow V^{\prime}_{6}$ such that $\mathsf{\Lambda}^{3}\varphi(A)=A^{\prime}$.
The double cover construction has been generalized in [17] to other EPW varieties $Y^{\geq\ell}_{A}$ and the double EPW surface defined below is in particular interesting.
Theorem 2.11 (Debarre–Kuznetsov).
Consider a Lagrangian subspace $A\subset\mathsf{\Lambda}^{3}V_{6}$ which does not contain decomposable elements and the associated EPW surface $Y^{\geq 2}_{A}\subset\mathbb{P}(V_{6})$, then there exists a unique double covering $f_{A}\colon\tilde{Y}^{\geq 2}_{A}\rightarrow Y_{A}^{\geq 2}$ branched over $Y_{A}^{\geq 2}$ such that
  • (a)
    the variety $\tilde{Y}_{A}^{\geq 1}$ is simply connected;
  • (b)
    the singular locus of is the finite set $f^{-1}_{A}(Y_{A}^{\geq 3})$;
  • (c)
    the variety is a $K3[2]$-type hyperkähler manifold once $Y_{A}^{\geq 3}=\varnothing$.
The variety $\tilde{Y}_{A}^{\geq 2}$ is called the double EPW surface associated with the Lagrangian subspace $A$.
The annihilator $A^{\perp}\subset\mathsf{\Lambda}^{3}V_{6}^{\vee}$ of a given Lagrangian subspace $A\subset\mathsf{\Lambda}^{3}V_{6}$ is also a Lagrangian subspace called the orthogonal of $A$. It contains no decomposable elements if and only if $A$ contains no decomposable elements. So one can define the dual EPW sextic $Y_{A^{\perp}}^{\geq 1}$, the double dual EPW sextic $\tilde{Y}_{A^{\perp}}^{\geq 1}$, the dual EPW surface $Y_{A^{\perp}}^{\geq 2}$, and the double dual EPW surface $\tilde{Y}_{A^{\perp}}^{\geq 2}$ for a Lagrangian subspace $A\subset\mathsf{\Lambda}^{3}V_{6}$ as the respective EPW varieties associated with $A^{\perp}$.
2.3. The period of Gushel–Mukai threefolds
Here let us restrict to the case of Gushel–Mukai threefolds. A Gushel–Mukai threefold $(X,H)$ determines a Lagrangian subspace $A$ in $\mathsf{\Lambda}^{3}V_{6}$ for the vector space $V_{6}=H^{0}(\mathbb{P}(W),\mathscr{I}_{X}(2))$. Together with the vector space $V_{5}$ in the definition of a Gushel–Mukai variety, one gets a triple $(V_{6},V_{5},A)$ which in turn determines $X$.
Definition 2.12.
A smooth Lagrangian data set is a collection $(V_{6},V_{5},A)$, where
  • $V_{6}$ is a $6$-dimensional vector space;
  • $V_{5}\subset V_{6}$ is a hyperplane;
  • $A\subset\mathsf{\Lambda}^{3}V_{6}$ is a Lagrangian subspace with no decomposable elements.
An isomorphism of smooth Lagrangian data sets between $(V_{6},V_{5},A)$ and $(V^{\prime}_{6},V^{\prime}_{5},A^{\prime})$ is a linear isomorphism $\varphi\colon V_{6}\rightarrow V^{\prime}_{6}$ such that $\varphi(V_{5})=V^{\prime}_{5}$ and $(\mathsf{\Lambda}^{3}\varphi)(A)=A^{\prime}$.
One has $\dim(A\cap\mathsf{\Lambda}^{3}V_{5})=3$ for the smooth Lagrangian data set $(V_{6},V_{5},A)$ associated with a special Gushel–Mukai threefold, and has $\dim(A\cap\mathsf{\Lambda}^{3}V_{5})=2$ for the smooth Lagrangian data set $(V_{6},V_{5},A)$ associated with an ordinary Gushel–Mukai threefold. Conversely, one can see
Theorem 2.13 (Debarre–Kuznetsov).
The map sending a Gushel–Mukai threefold to its smooth Lagrangian data induces a bijection between the set of Gushel–Mukai threefolds, and the set of smooth Lagrangian data $(V_{6},V_{5},A)$ such that $\dim(A\cap\mathsf{\Lambda}^{3}V_{5})\in\{2,3\}$.
Remark 2.14.
The quadric determining a Gushel–Mukai variety is not unique. In fact, every quadric contained in $\mathbb{P}(V_{6})-\mathbb{P}(V_{5})$ for a Gushel–Mukai variety $X$ gives $X$.
Since the double EPW sextic $\tilde{Y}_{A}^{\geq 1}$ is determined by the data $(V_{6},A)$, one cannot distinguish the Gushel–Mukai threefolds by their associated double EPW sextics directly.
Definition 2.15.
Two Gushel–Mukai threefolds are called period partner when they admit isomorphic double EPW sextics. Two Gushel–Mukai threefolds are called period dual when one associated double EPW sextics is the double dual EPW sextics for the other one.
Due to [7, Theorem 1.9], information of the period partners or duals for a Gushel–Mukai threefolds $X$ can be read from the Kuznetsov component ${\textbf{{Ku}}}(X)$. Once an addition data is included (serves as $V_{5}$), one can recover $X$ from ${\textbf{{Ku}}}(X)$ according to [29].
3. Stability conditions, moduli spaces, and on K3 surfaces
3.1. Stability conditions on triangulated categories
To reconstruct EPW varieties, we need to use the notion of stability conditions introduced by Bridgeland [10]. Here we will adapt a more recent formulation of stability conditions proposed in [6, Appendix 1] based on the locally-finite full stability conditions in [10, 11].
Here one fixes a triangulated category $\mathcal{D}$ and a finite rank lattice $\Lambda$ with a surjective homomorphism $\textup{{v}}\colon K(\mathcal{D})\twoheadrightarrow\Lambda$.
Definition 3.1.
A stability condition on $\mathcal{D}$ is a pair $\sigma=(\mathcal{P},Z)$ where
  • the central charge $Z$ is a linear map $Z\colon\Lambda\rightarrow\mathbb{C}$, and
  • the slicing $\mathcal{P}$ is a collection of full subcategories $\mathcal{P}(\phi)\subset\mathcal{D}$ for all $\phi\in\mathbb{R}$,
such that
  • $\mathcal{P}(\phi+1)=\mathcal{P}(\phi)[1]$ and $\operatorname{Hom}(\mathcal{P}(\phi_{1}),\mathcal{P}(\phi_{2}))=0$ for any $\phi_{1}>\phi_{2}$;
  • one has $Z(\textup{{v}}(E))\in\mathbb{R}_{>0}\cdot\exp(\pi\phi\sqrt{-1})$ for any $0\neq A\in\mathcal{P}(\phi)$;
  • every object $E$ of $\mathcal{D}$ admits a Harder–Narasimhan filtration
[…]
    with $A_{i}\in\mathcal{P}(\phi_{i})$ for all $i$.
  • there exists a quadratic form $Q$ on the vector space $\Lambda_{\mathbb{R}}$ such that $\ker(Z_{\mathbb{R}})$ is negative definite with respect to $Q$, and one has $Q(\textup{{v}}(E))\geq 0$ for any $A\in\mathcal{P}(\phi)$ and any $\phi\in\mathbb{R}$.
In this case, each $\mathcal{P}(\phi)$ is an abelian category. A non-zero object of $\mathcal{D}$ is called semistable with phase $\phi$ once it belongs to $\mathcal{P}(\phi)$ and is called stable once it is semistable and simple in $\mathcal{P}(\phi)$.
Remark 3.2.
Let $\sigma$ be stability condition, then the Harder–Narasimhan filtration for an object $E$ is unique up to isomorphism. In this case, the objects $A_{i}$ are called the factors of $E$. Two objects are called $S$-equivalence once they admitting the same Harder–Narasimhan factors.
Remark 3.3.
Let $\sigma=(\mathcal{P},Z)$ be stability condition, then any abelian category $\mathcal{P}(\phi)$ has finite length (see for example [10, page 330]) and an object in $\mathcal{P}(\phi)$ therefore admits a Jordan–Hölder filtration with uniquely determined Jordan–Hölder factors.
Remark 3.4.
Let $\sigma=(\mathcal{P},Z)$ be a stability condition on $\mathcal{D}$, then the smallest extension-closed subcategory $\mathcal{P}((0,1])\subset\mathcal{D}$ containing all the abelian subcategories $\mathcal{P}(\phi)$ for $\phi\in(0,1]$ is the heart of a bounded $t$-structure on $\mathcal{D}$. In this case, it is called the heart of $\sigma$.
This article will consider only certain stability conditions on triangulated categories arising from geometry.
Definition 3.5.
A triangulated category $\mathcal{D}$ is called geometric once it is equivalent to a triangulated subcategory of the bounded derived category of a smooth projective variety $X$ such that the embedding $\mathcal{D}\hookrightarrow{\textbf{{D}}}^{b}(X)$ has both the left and the right adjoint functors.
The Euler form $\chi$ on a smooth projective variety $X$ restricts to a geometric triangulated category $\mathcal{D}\subset{\textbf{{D}}}^{b}(X)$ and one defines $\textsf{{N}}(\mathcal{D}):=K(\mathcal{D})/\ker(\chi)$.
Definition 3.6.
A stability condition on a geometric triangulated category is called numerical once the surjection v is the quotient $K(\mathcal{D})\twoheadrightarrow\textsf{{N}}(\mathcal{D}):=K(\mathcal{D})/\ker(\chi)$.
Henceforth, a triangulated category $\mathcal{D}$ is always geometric and a stability condition is always numerical. According to [10], the set of (numerical) stability conditions on $\mathcal{D}$, if non-empty, admits a complex manifold structure $\operatorname{Stab}(\mathcal{D})$ with dimension equal to the rank of $\textsf{{N}}(\mathcal{D})$.
Also, there are two mutually-commutative group actions on the complex manifold $\operatorname{Stab}(\mathcal{D})$ due to [10, Lemma 8.2] as following.
The group $\operatorname{Aut}(\mathcal{D})$ acts on $\operatorname{Stab}(\mathcal{D})$ on the left, via the mapping
$\Phi.(\mathcal{P},Z):=(\mathcal{P}^{\prime},Z\circ\Phi_{*}^{-1})$
where $\Phi_{*}$ is the induced automorphism on $\textsf{{N}}(\mathcal{D})$ and $\mathcal{P}^{\prime}$ is defined by $\mathcal{P}^{\prime}(\phi):=\Phi(\mathcal{P}(\phi))$.
The universal cover
[…]
of the group $\operatorname{GL^{+}}(2,\mathbb{R})=\{M\in\operatorname{GL}(2,\mathbb{R})\,|\,\det(M)>0\}$ acts on the right on $\operatorname{Stab}(\mathcal{C})$, via
$(\mathcal{P}=\{\mathcal{P}(\phi)\}_{\phi\in\mathbb{R}},Z).(M,f):=(\mathcal{P}^{\prime}=\{\mathcal{P}(f(\phi))\}_{\phi\in\mathbb{R}},M^{-1}\circ Z).$
where we identify $\mathbb{C}$ with $\mathbb{R}\oplus\mathbb{R}\sqrt{-1}$ to make sense the composite $M^{-1}\circ Z\colon\textsf{{N}}(\mathcal{D})\rightarrow\mathbb{C}$.
3.2. Moduli spaces of semistable objects
Let $\sigma$ be a stability condition on a triangulated category $\mathcal{D}\subset{\textbf{{D}}}^{b}(X)$ and $v$ be an element in $\textsf{{N}}(\mathcal{D})$, then one can define a moduli stack $\mathfrak{M}_{\sigma}(v)$ of $\sigma$-semistable objects with numerical class $v$. A comprehensive reference is [2].
Definition 3.7.
The moduli stack $\mathfrak{M}_{\sigma}(v)\colon({\textbf{{Sch}}}/\mathbb{C})^{{\textbf{{op}}}}\rightarrow{\textbf{{Gpds}}}$ is defined by setting
[…]
for any locally finitely generated complex scheme $T$. Similarly, one defines $\mathfrak{M}^{\textsf{{st}}}_{\sigma}(v)$.
A stability condition is called proper once, for any $v$, the moduli stack $\mathfrak{M}_{\sigma}(v)$ is an algebraic stack of finite type, and there is an open immersion $\mathfrak{M}_{\sigma}^{\textsf{{st}}}(v)\subset\mathfrak{M}_{\sigma}(v)$, and $\mathfrak{M}_{\sigma}(v)$ admits a proper good moduli space $\textsf{{M}}_{\sigma}(v)$ in the sense of [1], and the open substack $\mathfrak{M}_{\sigma}^{\textsf{{st}}}(v)$ is a $\mathbb{G}_{m}$-gerbe over the counterpart $\textsf{{M}}^{\textsf{{st}}}_{\sigma}(v)\subset\textsf{{M}}_{\sigma}(v)$. All the stability conditions in this article are proper.
Such a good moduli space $\textsf{{M}}_{\sigma}(v)$ is an algebraic space and the (closed) points in $\textsf{{M}}_{\sigma}(v)$ are $S$-equivalent classes of semistable objects with phase ranging in (1,1](-1,1].
Moreover, the moduli space $\textsf{{M}}_{\sigma}(v)$ varies with respect to the variation of the stability condition $\sigma$ in the complex manifold $\operatorname{Stab}(\mathcal{D})$.
Proposition 3.8 ([11, Section 9] and [4, Proposition 2.3]).
Given $v\in\textsf{{N}}(\mathcal{D})$, then there exists a locally finite set of walls (real codimension one submanifolds with boundary) in the complex manifold $\operatorname{Stab}(\mathcal{D})$, depending only on $v$, such that:
  • The sets of $\sigma$-semistable and $\sigma$-stable objects of class $v$ remain unchanged for any stability condition $\sigma$ in the same chamber;
  • When $\sigma$ lies on a single wall in $\operatorname{Stab}(\mathcal{D})$, then there is a $\sigma$-semistable object that is unstable in one of the adjacent chambers, and semistable in the other adjacent chamber.
Suppose in addition that $v$ is primitive in the lattice $\textsf{{N}}(\mathcal{D})$, then a stability condition $\sigma$ lies on a wall with respect to $v$ if and only if there exists a strictly $\sigma$-semistable object of class $v$.
In general, the classification of walls are not available but one always has the following two situations for a stability condition.
Definition 3.9.
Let $v$ be a class in $\textsf{{N}}(\mathcal{D})$, a stability condition $\sigma$ on a triangulated category $\mathcal{D}$ is called $v$-generic when it is not contained in any wall with respect to $v$.
Definition 3.10.
Let $W\subset\operatorname{Stab}(\mathcal{D})$ be a wall with respect to a given class $v\in\textsf{{N}}(\mathcal{D})$ such that one has $\textsf{{M}}^{\textsf{{st}}}_{\sigma}(v)\neq\varnothing$ for some $\sigma\in\operatorname{Stab}(\mathcal{D})$ close to $W$, then $W$ is called a totally semistable wall if one has $\textsf{{M}}^{\textsf{{st}}}_{\sigma_{0}}(v)=\varnothing$ for any $\sigma_{0}\in W$ which is not in any other wall.
The walls and wall-crossing behaviors of the moduli space $\textsf{{M}}_{\sigma}(v)$ are only well-understood for some stability conditions on K3 surfaces [4, 5, 35] and Enriques surfaces [36]. Since a strongly smooth Gushel–Mukai surface is in particular a K3 surface, the article uses some results regarding walls and wall-crossing for K3 surfaces which are recollected in what follows.
3.3. Some stability conditions on K3 surfaces
A connected component of the complex manifold $\operatorname{Stab}({\textbf{{D}}}^{b}(S))$ for a K3 surface $S$ is introduced in [11]. In this case, the Mukai vector
$v(E)=\textup{{ch}}(E).\sqrt{\textup{{td}}(S)}=(\operatorname{rank}(E),c_{1}(E),\operatorname{rank}(E)+c_{1}(E)^{2}/2-c_{2}(E))$
identifies the numerical Grothendieck group $\textsf{{N}}(S):=\textsf{{N}}({\textbf{{D}}}^{b}(S))$ with the algebraic Mukai lattice
$\textsf{{H}}(S,\mathbb{Z})=H^{0}(S,\mathbb{Z})\oplus\operatorname{NS}(S)\oplus H^{4}(S,\mathbb{Z})$
where $\operatorname{NS}(S)$ is the Néron–Severi group of the K3 surface. The lattice $\textsf{{H}}(S,\mathbb{Z})$ is endowed with the Mukai pairing $\langle(r_{1},\Delta_{1},s_{1}),(r_{2},\Delta_{2},s_{2})\rangle=\Delta_{1}.\Delta_{2}-r_{1}.s_{2}-r_{2}.s_{1}$ where $\Delta_{1}\cdot\Delta_{2}$ is the usual intersection product in $H^{2}(S,\mathbb{Z})$ and carries a canonical weight two Hodge structure from the one on the integer cohomology group $H^{0}(S,\mathbb{Z})\oplus H^{2}(S,\mathbb{Z})\oplus H^{4}(S,\mathbb{Z})$.
An ample class $\omega\in\operatorname{NS}(S)_{\mathbb{R}}$ defines a notion of $\mu_{\omega}$-slope stability on ${\textbf{{Coh}}}(S)$, by declaring the slope $\mu_{\omega}(\mathscr{F}):=c_{1}(\mathscr{F})\cdot\omega/\operatorname{rank}(\mathscr{F})$ for a torsion-free sheaf $\mathscr{F}$ and $\mu_{\omega}(\mathscr{F})=+\infty$ for a torsion $\mathscr{F}$. Truncating the HN filtrations at the number $\beta\cdot\omega$, a pair $(\omega,\beta)\in\operatorname{NS}(S)^{2}_{\mathbb{R}}$ with $\omega$ ample determines a torsion pair $(\mathcal{T}_{\omega,\beta},\mathcal{F}_{\omega,\beta})$ on ${\textbf{{Coh}}}(S)$ such that
[…]
This torsion pair gives a bounded $t$-structure on ${\textbf{{D}}}^{b}(S)$ with heart
$\mathcal{A}_{\omega,\beta}=\{E\in{\textbf{{D}}}^{b}(S)\,|\,H^{i}(E)=0\textup{ for }i\notin\{-1,0\},H^{-1}(E)\in\mathcal{F}_{\omega,\beta},H^{0}(E)\in\mathcal{T}_{\omega,\beta}\}$
according to [22], on which one can define a stability condition $\sigma_{\omega,\beta}$ under certain conditions.
Definition 3.11.
An object $E$ in ${\textbf{{D}}}^{b}(S)$ is called spherical once $\textup{{R}}\operatorname{Hom}(E,E)=\mathbb{C}[0]\oplus\mathbb{C}[-2]$.
Proposition 3.12 ([11, Lemma 6.2 and Section 7]).
Consider a K3 surface $S$ and two classes $\omega,\beta\in\operatorname{NS}(S)_{\mathbb{R}}$ with $\omega$ ample, then there exists a stability condition $\sigma_{\omega,\beta}$ with central charge
$Z_{\omega,\beta}\colon\textsf{{H}}(S,\mathbb{Z})\rightarrow\mathbb{C},\quad(r,\Delta,s)\mapsto\frac{1}{2}(2\beta\cdot\Delta-2s+r(\omega^{2}-\beta^{2}))+(\Delta-r\beta)\cdot\omega\sqrt{-1}$
and heart $\mathcal{A}_{\omega,\beta}$ if and only if $Z_{\omega,\beta}(\mathscr{F})\notin\mathbb{R}_{\leq 0}$ for any spherical sheaf $\mathscr{F}$ on $S$.
Example 3.13.
Let $(S,H)$ be a strongly smooth Gushel–Mukai surface endowed with the canonical degree $10$ polarization, then one can check that $(\omega,\beta)=(1/5H,-2/5H)$ defines a stability condition $\sigma_{S}:=\sigma_{\omega,\beta}$ on ${\textbf{{D}}}^{b}(S)$ with central charge
$Z_{\sigma_{S}}(r,\Delta,s)=-\left(\frac{3}{5}r+\frac{2}{5}H\cdot\Delta+s\right)+\left(\frac{4}{5}r+\frac{1}{5}\Delta\cdot H\right)\sqrt{-1}$
using the above criterion and Proposition 2.4.
The stability conditions $\sigma_{\omega,\beta}$ on a K3 surface $S$ are proper and contained in a connected component $\operatorname{Stab}^{{\dagger}}(S)$ of $\operatorname{Stab}({\textbf{{D}}}^{b}(S))$ according to [11, 45]. Moreover, the following statement asserts that, for a certain vector $v\in\textsf{{H}}(S,\mathbb{Z})$ and a given polarization $H$ on $S$, the moduli space $\textsf{{M}}_{\sigma}(v)$ is equal to the moduli space $\textsf{{M}}_{H}(v)$ of Gieseker semistable sheaves on $(S,H)$ with Mukai vector $v$ once the stability condition $\sigma$ belongs to a special chamber in $\operatorname{Stab}^{{\dagger}}(S)$
Proposition 3.14 ([4, 11, 45]).
Consider a polarized K3 surface $(S,H)$ and a stability condition $\sigma_{\omega,\beta}$ on ${\textbf{{D}}}^{b}(S)$ such that $\omega,\beta\in\mathbb{R}[H]$, then for any given $v\in\textsf{{H}}(S,\mathbb{Z})$ with $\textsf{{M}}_{H}(v)\neq\varnothing$ there exists a real number $t_{0}=t(v)$ such that $\textsf{{M}}_{\sigma_{t\omega,\beta}}(v)=\textsf{{M}}_{H}(v)$ for any $t\geq t_{0}$.
In general, the possible walls in $\operatorname{Stab}^{{\dagger}}(S)$ with respect to a given primitive Mukai vector are classified in [4] and the criteria to determine a wall can be find in [5]. These results will be recalled in the subsequent subsection.
3.4. The walls and wall-crossing on K3 surfaces
Let $S$ be a K3 surface and $v\in\textsf{{H}}(S,\mathbb{Z})$ be a primitive Mukai vector with $\langle v,v\rangle\geq-2$. Suppose that $W\subset\operatorname{Stab}^{{\dagger}}(S)$ be a wall respect to $v$ in the sense of Proposition 3.8 and $\sigma_{0}$ is a generic point on $W$, then for any two $v$-generic stability conditions $\sigma_{\pm}$ in the adjoint chambers with respect to $W$ one has a birational map
$\textsf{{M}}_{\sigma_{-}}(v)\dashrightarrow\textsf{{M}}_{\sigma_{+}}(v)$
and two contractions $\textsf{{M}}_{\sigma_{\pm}}(v)\rightarrow\bar{M}_{\pm}$ contracting $S$-equivalent objects by [4, Theorem 1.4].
To determine a potential wall $W$, one introduces a rank two primitive sublattice $\textbf{{H}}_{W}$ of the Mukai lattice $\textsf{{H}}(S,\mathbb{Z})$ such that $w\in\textbf{{H}}_{W}$ if and only if $\Im(Z(w)/Z(v))=0$ for any $\sigma\in W$.
Proposition 3.15 ([5, Proposition 5.1]).
Under the setups in this subsection, the Mukai vector of any HN factor of an object $A$ in $\textsf{{M}}_{\sigma_{+}}(v)$ with respect to $\sigma_{-}$ is contained in $\textbf{{H}}_{W}$, and vice versa. The same holds for any object in $\textsf{{M}}_{\sigma_{0}}(v)$. Similarly, any Jordan–Hölder factor of an object in $A\in\textsf{{M}}^{\textsf{{st}}}_{\sigma_{0}}(w)$ for some $w\in\textbf{{H}}_{W}$ will have Mukai vector contained in $\textbf{{H}}_{W}$.
In this case, using the existence of certain semistable objects, one can determine the type of the wall $W$ according to specific vectors in $\textbf{{H}}_{W}$ by [5, Theorem 5.7] when $\langle v,v\rangle>0$.
Definition 3.16.
Let $\langle v,v\rangle>0$, then the wall $W$ is called a fake wall once there are no curves in $\textsf{{M}}_{\sigma_{\pm}}(v)$ that are $S$-equivalent to each other respect to $\sigma_{0}$; a flopping wall once $\bar{M}_{+}=\bar{M}_{-}$ and the induced birational map $\textsf{{M}}_{\sigma_{-}}(v)\dashrightarrow\textsf{{M}}_{\sigma_{+}}(v)$ induces a flopping contraction; a divisorial wall once the induced contractions $\textsf{{M}}_{\sigma_{\pm}}(v)\rightarrow\bar{M}_{\pm}$ are divisorial contractions.
Theorem 3.17.
Under the setups in this subsection and assume that $\langle v,v\rangle>0$, then the potential wall $W$ is a totally semistable wall if and only if there exists either a vector $w\in\textbf{{H}}_{W}$ such that $\langle w,w\rangle=0$ and $\langle v,w\rangle=1$, or such that $\langle w,w\rangle=-2$, $\Re(Z(w)/Z(v))>0$ and $\langle v,w\rangle<0$.
In addition, $W$ is a wall inducing a divisorial contraction if and only if there exists a vector $w$ in $\textbf{{H}}_{W}$ such that either $\langle w,w\rangle=0$ and $\langle v,w\rangle\in\{1,2\}$ or $\langle w,w\rangle=-2$ and $\langle v,w\rangle=0$.
Otherwise $W$ is either a flopping wall with respect to $v$ or not a wall. In this article, only vector $v$ with $\langle v,v\rangle=2$ will be involved, so one has the following special criterion.
Theorem 3.18.
Suppose that $W$ is not a totally semistable wall or a divisorial wall and the vector $v$ satisfies $\langle v,v\rangle=2$, then $W$ is a flop wall if and only if there exists a vector $w\in\textbf{{H}}_{W}$ such that $\langle w,w\rangle=-2$ and $\langle v,w\rangle=1$.
In this case, according to [5, Proposition 9.4 and Section 14], the Jordan–Hölder factors for a strictly semistable object $A$ in $\textsf{{M}}_{\sigma_{0}}(v)$ are spherical $\sigma_{\pm}$-stable objects. Moreover, one notices that the vector $w^{\prime}=v-w$ also satisfies $\langle w^{\prime},w^{\prime}\rangle=-2$ and $\langle v,w^{\prime}\rangle=1$ in the setting of Theorem 3.18, so a general flopping wall is induced by two spherical stable objects.
To identify the spherical factors for such a flopping wall, one also needs to understand the walls for $\textsf{{M}}_{\sigma_{0}}(v)$ with respect to spherical classes.
Proposition 3.19 ([9, Proposition 4.8]).
Consider a spherical class $v$ and $A\in\textsf{{M}}_{\sigma_{+}}(v)$ which is not $\sigma_{0}$-stable, then the set of Jordan-Holder factors for $A$ contains only two spherical objects.
4. The passage to special Gushel–Mukai threefolds
4.1. Group actions and equivariant categories
We fix a finite group $G$ and a category $\mathcal{D}$.
Definition 4.1.
A (right) action of $G$ on $\mathcal{D}$ is an assignment ${\textup{{{A}}}}_{(-)}\colon G\rightarrow\operatorname{Aut}(\mathcal{D})$ and a set of natural transformations $\{\epsilon_{g,h}\colon{\textup{{{A}}}}_{gh}\stackrel{{\scriptstyle\sim}}{{\rightarrow}}{\textup{{{A}}}}_{h}\circ{\textup{{{A}}}}_{g}\,|\,g,h\in G\}$ such that
${{\textup{{{A}}}}_{g_{1}g_{2}g_{3}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\epsilon_{g_{1},g_{2}g_{3}}}$$\scriptstyle{\epsilon_{g_{1}g_{2},g_{3}}}$${{\textup{{{A}}}}_{g_{3}}\circ{\textup{{{A}}}}_{g_{1}g_{2}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\textup{{{A}}}}_{g_{3}}\epsilon_{g_{1},g_{2}}}$${{\textup{{{A}}}}_{g_{2}g_{3}}\circ{\textup{{{A}}}}_{g_{1}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\epsilon_{g_{2},g_{3}}{\textup{{{A}}}}_{g_{1}}}$${{\textup{{{A}}}}_{g_{3}}\circ{\textup{{{A}}}}_{g_{2}}\circ{\textup{{{A}}}}_{g_{1}}}$
commutes for all $g_{1},g_{2},g_{3}\in G$.
Given an action of $G$ on $\mathcal{D}$, one can always construct the equivariant category $\mathcal{D}_{G}$ with respect to the $G$-action on $\mathcal{D}$ whose objects and morphisms are respectively the $G$-equivariant objects and $G$-equivariant morphisms.
Definition 4.2.
A $G$-equivariant object of $\mathcal{D}$ is a pair $(A,\phi)$, where $A$ is an object of $\mathcal{D}$ and $\phi=\{\phi_{g}\}_{g\in G}$ is a collection of isomorphisms $\phi_{g}:A\xrightarrow{\sim}{\textup{{{A}}}}_{g}A$ for $g\in G$, such that the diagram
[…]
commutes for all $g,h\in G$. The equivariant structure $\phi$ is often omitted from the notation.
Definition 4.3.
A $G$-equivariant morphism $(A,\phi^{A})\rightarrow(B,\phi^{B})$ of $G$-equivariant objects of $\mathcal{D}$ is a morphism $f\colon A\rightarrow B$ in $\mathcal{D}$ such that the diagram
[…]
commutes for every $g\in G$.
Suppose now that $G$ is abelian, then one can define a natural action of the dual group $G^{\vee}:=\operatorname{Hom}(G,\mathbb{C}^{\times})$ on $\mathcal{D}_{G}$ from a $G$-action on $\mathcal{D}$: for $\chi\in G^{\vee}$ and $(A,\phi)\in\mathcal{D}$ one sets
$\psi_{\chi}(A,\{\phi_{g}\}_{g\in G})=(A,\{\phi_{g}\}_{g\in G})\otimes\chi=(A,\{\phi_{g}\cdot\chi(g)\})$
The equivariant objects $\psi_{\chi_{1}}(\psi_{\chi_{2}}(A,\phi))$ and $\psi_{\chi_{2}\chi_{1}}(A,\phi)$ are the same, so one can just let the required natural transformations $\psi_{\chi_{1}}\circ\psi_{\chi_{2}}\rightarrow\psi_{\chi_{2}\chi_{1}}$ be identities.
Theorem 4.4 ([19, Theorem 4.2]).
Consider a finite abelian group $G$ acting on an idempotent complete category $\mathcal{D}$, then there exists an equivalence
${\textup{{{X}}}}\colon(\mathcal{D}_{G})_{G^{\vee}}\cong\mathcal{D}$
such that the $G$-action is identified with the $(G^{\vee})^{\vee}$-action by $(G^{\vee})^{\vee}\cong G$.
There are mutually left and right adjoint functors: the forgetful functor
${\textup{{{Forg}}}}\colon\mathcal{D}_{G}\rightarrow\mathcal{D}$
defined by forgetting the $G$-equivariant structure, and the inflation functor
${\textup{{{Inf}}}}\colon\mathcal{D}\rightarrow\mathcal{D}_{G}$
given on objects by sending $A$ to $\bigoplus_{g\in G}{\textup{{{A}}}}_{g}A$ with the natural $G$-equivariant structure (one can see, for example, [19, Lemma 3.8]). Moreover, one has the following observation:
Proposition 4.5.
Under the assumptions of Theorem 4.4, there is an equivalence of functors
${\textup{{{Forg}}}}\cong{\textup{{{Inf}}}}\circ{\textup{{{X}}}}\colon(\mathcal{D}_{G})_{G^{\vee}}\rightarrow\mathcal{D}_{G}$
where Forg is the one for the induced $G^{\vee}$-action on $\mathcal{D}_{G}$.
Suppose in addition that both $\mathcal{D}$ and $\mathcal{D}_{G}$ are triangulated categories, then $(\mathcal{D}_{G})_{G^{\vee}}$ comes with a natural class of exact triangles, consisting of those triangles whose image under
${\textup{{{Forg}}}}\colon(\mathcal{D}_{G})_{G^{\vee}}\rightarrow\mathcal{D}_{G}$
are exact. Then, due to Proposition 4.5, the equivalence X in Theorem 4.4 respects exact triangles and therefore an equivalence of triangulated categories.
4.2. Equivariant stability conditions and covering maps
Let $G$ be a finite group action on a triangulated category $\mathcal{D}$. Then one can construct new stability conditions on $\mathcal{D}_{G}$ out of the stability conditions on $\mathcal{D}$ under suitable compatible conditions. To avoid unnecessary technical notions, we will only consider the case where $\mathcal{D}={\textbf{{D}}}^{b}(X)$ for a smooth projective variety $X$ and the group $G$ is the cycle group $\mathfrak{S}_{2}$. One consults [43] for more general statements.
Theorem 4.6.
[34, 44, 43] Suppose that the involution $\Pi$ on ${\textbf{{D}}}^{b}(X)$ is a generator of the group action $\mathfrak{S}_{2}$ on $\mathcal{D}={\textbf{{D}}}^{b}(X)$, and let $\sigma=(\mathcal{P},Z)$ be a stability condition on $\mathcal{D}$ such that for each $\phi\in\mathbb{R}$ one has $\Pi(\mathcal{P}(\phi))=\mathcal{P}(\phi)$ and $Z\circ\Pi_{*}=Z$ where $\Pi_{*}$ is the action of $\Pi$ on $\textsf{{N}}(X)$, then
$\mathcal{P}_{\mathfrak{S}_{2}}(\phi)$ $:=\{A\in\mathcal{D}_{\mathfrak{S}_{2}}\,|\,{\textup{{{Forg}}}}(A)\in\mathcal{P}(\phi)\}$
$Z_{\mathfrak{S}_{2}}$ $=Z\circ\textup{{forg}}\colon\textsf{{N}}(\mathcal{D}_{\mathfrak{S}_{2}})\rightarrow\mathbb{C}$
determines a stability condition $\sigma_{\mathfrak{S}_{2}}$ on $\mathcal{D}_{\mathfrak{S}_{2}}$.
In the construction from Theorem 4.6, both the central charge $Z_{\mathfrak{S}_{2}}$ and the slicing $\mathcal{P}_{\mathfrak{S}_{2}}$ of the stability condition $\sigma_{\mathfrak{S}_{2}}$ are preserved by the dual group $\mathfrak{S}_{2}^{\vee}$ action on $\mathcal{D}_{\mathfrak{S}_{2}}$. So we may apply the theorem again to obtain a stability condition $(\sigma_{\mathfrak{S}_{2}})_{\mathfrak{S}_{2}^{\vee}}$ on $(\mathcal{D}_{\mathfrak{S}_{2}})_{\mathfrak{S}_{2}^{\vee}}$. This iterated construction is compatible with the equivalence in Theorem 4.4, up to a multiplication by $2$.
Proposition 4.7.
In the setup of Theorem 4.6, the stability condition $(\sigma_{\mathfrak{S}_{2}})_{\mathfrak{S}_{2}^{\vee}}$ corresponds to the stability condition $2\cdot\sigma=(\mathcal{P},2Z)$ under the equivalence ${\textup{{{X}}}}\colon(\mathcal{D}_{\mathfrak{S}_{2}})_{\mathfrak{S}_{2}^{\vee}}\cong\mathcal{D}$ of Theorem $4.$
4.3. The Kuznetsov components and the involutions
Let $X$ be a special Gushel–Mukai threefold endowed with the canonical degree $10$ ample polarization $\mathscr{O}_{X}(1)$, and let $\mathscr{U}_{X}$ be the pullback of the tautological bundle on $\operatorname{Gr}(2,V_{5})$ along the morphism $X\rightarrow\operatorname{Gr}(2,V_{5})$ in Remark 2.6, then ${\textbf{{D}}}^{b}(X)$ has a geometric triangulated subcategory
${\textbf{{Ku}}}(X):=\{E\in{\textbf{{D}}}^{b}(X)\,|\,\textsf{R}\operatorname{Hom}(\mathscr{O}_{X},E)=\textsf{R}\operatorname{Hom}(\mathscr{U}^{\vee}_{X},E)=0\}$
which is called the Kuznetsov component of $X$.
The numerical Grothendieck group $\textsf{{N}}({\textbf{{Ku}}}(X))$ has rank 2. Denotes the first Chern class of the polarization $\mathscr{O}_{X}(1)$ by $h$, then one has a basis $\kappa_{1},\kappa_{2}$ of $\textsf{{N}}({\textbf{{Ku}}}(X))$ with
$\textup{{ch}}(\kappa_{1})=1-\frac{1}{5}h^{2}\quad\textup{and}\quad\textup{{ch}}(\kappa_{1})=2-h+\frac{1}{12}h^{3}$
such that the Euler form is given in this basis by $I_{2}$. One consults [30, 43] for details.
The opposite of $X$ is a strongly smooth Gushel–Mukai surface $S$ and one has a covering involution $f\colon X\rightarrow X$ for the double cover $X\rightarrow M_{S}$ branched over $S\subset M_{S}$. The direct image of the covering involution restricts to an involution on the Kuznetsov component ${\textbf{{Ku}}}(X)$ and induces a $\mathfrak{S}_{2}$-action on ${\textbf{{Ku}}}(X)$. According to [30, 43], the equivariant category ${\textbf{{Ku}}}(X)_{\mathfrak{S}_{2}}$ is equivalent to ${\textbf{{D}}}^{b}(S)$ and the dual $\mathfrak{S}_{2}$-group action on the equivariant category is induced by an involution on ${\textbf{{D}}}^{b}(S)$ say $\Pi$. The involution $\Pi$ induces a Hodge isometry on $\textsf{{H}}(S,\mathbb{Z})$ whose invariant lattice is spanned by $v_{1}=(1,0,-1)$ and $v_{2}=(2,-H,2)$.
The forgetful functor ${\textup{{{Forg}}}}\colon{\textbf{{D}}}^{b}(S)\rightarrow{\textbf{{Ku}}}(X)$ induces a group homomorphism forg from the lattice $\textsf{{H}}(S,\mathbb{Z})$ to the lattice $\textsf{{N}}({\textbf{{Ku}}}(X))$ and the inflation functor ${\textup{{{Inf}}}}\colon{\textbf{{Ku}}}(X)\rightarrow{\textbf{{D}}}^{b}(S)$ induces a group homomorphism $\textup{{inf}}\colon\textsf{{N}}({\textbf{{Ku}}}(X))\rightarrow\textsf{{H}}(S,\mathbb{Z})$.
Proposition 4.8 ([33, Propisition 6.6]).
Under the setups above, one has
$\textup{{forg}}(r,\Delta,s)$ $=(s-r)\kappa_{1}+(2r+2s+\Delta\cdot H)\kappa_{2}$
$\textup{{inf}}(a\kappa_{1}+b\kappa_{2})$ $=-a(1,0,-1)-b(2,-H,2)$
for any vector $(r,\Delta,s)\in\textsf{{H}}(S,\mathbb{Z})$ and integers $a,b$.
5. Double duel EPW sextics as Bridgeland moduli spaces
In this section, we are going to show that
$\textsf{{M}}_{\sigma_{S}}(v_{1})\stackrel{{\scriptstyle\sim}}{{\rightarrow}}\tilde{Y}^{\geq 1}_{A(S)^{\perp}}$
for any strongly smooth Gushel–Mukai surface $S$.
5.1. Birational model of the double dual EPW sextics
To achieve the isomorphism, one needs the description of the associated double dual EPW sextic in [18, Section 7.5].
At first, one notices that each line $L\subset S$ gives a Lagrangian submanifold $\mathbb{P}^{2}\cong(\mathbb{P}^{1})^{[2]}$ contained in $S^{[2]}$. In this case, one has a Mukai flop
${\hat{X}_{S}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${S^{[2]}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g}$${X_{S}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\bar{X}_{S}}$
with respect to all these Lagrangian planes. The variety $\bar{X}_{S}$ is a subvariety of $\operatorname{Gr}(2,W)$ and the contraction $S^{[2]}\rightarrow\bar{X}_{S}$ is given by sending a point in $S^{[2]}$ to its spanned line in $\operatorname{Gr}(2,W)$.
Then one considers the Fano variety of lines $F(M_{S})\cong\mathbb{P}^{2}$ in the Grassmannian hull $M_{S}$ of the strongly smooth Gushel–Mukai surface $S$. It canonically contains the Fano variety of lines $F(S)$ on $S$, and embeds into the contracted space $\bar{X}_{S}$. So the blow-up $Z_{S}$ of $F(M_{S})$ along the set $F(S)$ of points embeds into $S^{[2]}$. The strict transformation of the blow-up $Z_{S}$ along the Mukai flop $g$ is isomorphic to $F(M_{S})$ and will be again denoted by $F(M_{S})$.
Also, one notices that a conic (including the degenerated ones) and a line in $S$ cannot meet exactly at a length $2$ subscheme due to Theorem 2.4. So the dual plane $\Sigma_{C}^{\vee}$ parameterizing lines in the plane $\Sigma_{C}$ spanned by $C$ can be realized as a subvariety in $X_{S}$.
Indeed, suppose that $C$ is smooth, then every line in $\Sigma_{C}$ intersects $C$ in a length $2$ subscheme which is not contained in the exceptional locus of $g$ as the conic $C$ cannot meet a line in $S$ at a subscheme of length $2$. So $\Sigma_{C}^{\vee}\cong\Sigma_{C}^{[2]}\subset S^{[2]}$ is not contained in the exceptional locus of $g$ and hence embeds into $X_{S}$. The singular case is a bit complicated. Suppose for example that $C$ is a union of two lines, then $\Sigma_{C}^{[2]}$ is the union of the two Lagrangian planes spanned by these lines and the blow-up of $\Sigma_{C}^{\vee}$ along the two points represented by the lines. So one can see that $\Sigma_{C}^{\vee}$ is the strict transformation of $\Sigma_{C}^{[2]}$ via the flop $g$ and hence embeds into $X_{S}$.  [^1] These details are not included in [18] and we thank Sasha Kuznetsov for the further explanation.
[…]
$\bm{p}_{S}\in Y^{\geq 3}_{A(S)^{\perp}}=\{[v]\in\mathbb{P}(V_{6}^{\vee})\,|\,\dim(A^{\perp}_{S}\cap\left(v\land\mathsf{\Lambda}^{2}V_{6}^{\vee}\right))\geq 3\}$
determined by the subspace $V_{5}\subset V_{6}$ according to [13, Theorem 3.10], and $Y^{\geq 3}_{A(S)^{\perp}}-\{\bm{p}_{S}\}$ is equal to the set of conics in $S$. The point $\bm{p}_{C}\in Y^{\geq 3}_{A(S)^{\perp}}$ corresponding to a conic $C\subset S$ is given by the unique 5-dimensional subspace $U_{5}\subset V_{6}$ determined by the linear span of $C$ in $\mathbb{P}(W)$.
Theorem 5.1 ([18, Theorem 7.7 and Remark 7.9]).
There is a diagram
$S^{[2]}\stackrel{{\scriptstyle g}}{{\dashrightarrow}}X_{S}\stackrel{{\scriptstyle f^{+}}}{{\longrightarrow}}\tilde{Y}_{A(S)^{\perp}}^{\geq 1}$
where the birational map $g$ is the Mukai flop of all Lagrangian planes in $S^{[2]}$ corresponding to lines in $S$ and the morphism $f^{+}$ is a symplectic resolution of singularities. In particular, the map $f^{+}$ contracts $F(M_{S})\subset X_{S}$ to the singular point $\tilde{\bm{p}}_{S}$ and, for any other point corresponding to a conic $C\subset S$, the dual projective plane $\Sigma^{\vee}_{C}$ to that point.
Remark 5.2.
The Mukai flop $g$ coincides with the flop described in [39, (4.2.4)]. Once $S$ does not contain lines and conics, the resolution $f^{+}$ coincides with the one in [39, (4.2.2)].
In the following subsections, we will recover diagram by studying the wall-crossing behavior of the moduli space $\textsf{{M}}_{\sigma_{t}}(v_{1})$ for the stability condition $\sigma_{t}:=\sigma_{\sqrt{t}/5H,-2/5H}$ with $t\geq 1$.
5.2. Exclusion of the non-flopping walls
In this subsection, we will show that either $\sigma_{t}$ is not on a wall with respect to $v_{1}$ or on a flopping wall with respect to $v_{1}$.
The central charge of the stability condition $\sigma_{t}$ is
$Z_{\sigma_{t}}(r,\Delta,s)=\left((\frac{t-4}{5})r-\frac{2}{5}H\cdot\Delta-s\right)+\left(\frac{4}{5}r+\frac{1}{5}\Delta\cdot H\right)\sqrt{-t}$
and the associated lattice $\textbf{{H}}_{W}$ of a potential wall $W$ for $v_{1}=(1,0,-1)$ is characterized by
$\Im\frac{Z_{\sigma_{t}}(r,\Delta,s)}{Z_{\sigma_{t}}(1,0,-1)}=\Im\left(\frac{((t-4)r-2\Delta\cdot H-5s)+(\Delta\cdot H+4r)\sqrt{-t}}{(t+1)+4\sqrt{-t}}\right)=0$
or equivalently the condition $20r+(t+9)\Delta\cdot H+20s=0$.
Proposition 5.3.
There is no class $w\in\textbf{{H}}_{W}$ such that $\langle w,w\rangle=-2$ and $\langle v,w\rangle=0$.
Proof.
Suppose otherwise that there such a vector $w=(r,\Delta,s)$, then one has has $(t+9)\Delta\cdot H=-40r$ and $\Delta^{2}=2r^{2}-2$. It follows
$20r^{2}-20=10\Delta^{2}\leq(\Delta\cdot H)^{2}=\left(\frac{40r}{t+9}\right)^{2}$
or equivalently $((t+9)^{2}-80)r^{2}\leq(t+9)^{2}$. So $r^{2}\leq 5$ and all the possible cases are
  • $r=0,t\geq 1$ and $\Delta^{2}=-2$ and $\Delta\cdot H=0$;
  • $r=\pm 1,t=1$ and $\Delta^{2}=0$ and $\Delta\cdot H=\pm 4$;
  • $r=\pm 1,t=11$ and $\Delta^{2}=0$ and $\Delta\cdot H=\pm 2$;
  • $r=\pm 1,t=31$ and $\Delta^{2}=0$ and $\Delta\cdot H=\pm 1$;
  • $r=\pm 2,t=1$ and $\Delta^{2}=6$ and $\Delta\cdot H=\pm 8$.
The first case is absurd as $|\Delta|$ should contain an effective class, and the other cases contradict to Proposition 2.4 (in the last case one has $(H\pm\Delta)^{2}=0$ and $(H\pm\Delta)\cdot H=2$). ∎
Proposition 5.4.
There is no vector $w\in\textbf{{H}}_{W}$ such that $\langle w,w\rangle=0$ and $\langle v,w\rangle=1$.
Proof.
Suppose otherwise that there exists such a vector $w=(r,\Delta,s)$, then one has $(t+9)\Delta\cdot H=-40r+20$ and $\Delta^{2}=2r^{2}-2r$. It follows
$20r^{2}-20r=10\Delta^{2}\leq(\Delta\cdot H)^{2}=\left(\frac{40r-20}{t+9}\right)^{2}$
or equivalently $((t+9)^{2}-80)r^{2}-((t+9)^{2}-80)r\leq 20$. So $r\in\{0,1\}$ and all possible cases are
  • $r=0,t=1$ and $\Delta^{2}=0$ and $\Delta\cdot H=2$;
  • $r=0,t=11$ and $\Delta^{2}=0$ and $\Delta\cdot H=1$;
  • $r=1,t=1$ and $\Delta^{2}=0$ and $\Delta\cdot H=-2$;
  • $r=1,t=11$ and $\Delta^{2}=0$ and $\Delta\cdot H=-1$;
which always conflict with Proposition 2.4. ∎
Proposition 5.5.
There is no vector $w\in\textbf{{H}}_{W}$ such that $\langle w,w\rangle=0$ and $\langle v,w\rangle=2$.
Proof.
Suppose otherwise that there exists such a vector $w=(r,\Delta,s)$, then one has $(t+9)\Delta\cdot H=-40r+40$ and $\Delta^{2}=2r^{2}-4r$. It follows
$20r^{2}-40r=10\Delta^{2}\leq(\Delta\cdot H)^{2}=\left(\frac{40r-40}{t+9}\right)^{2}$
or equivalently $((t+9)^{2}-80)r^{2}-2((t+9)^{2}-80)r\leq 80$. So $r\in\{-1,0,1,2,3\}$ and one can check as in Proposition 5.3 and Proposition 5.4 that all possible cases are invalid. ∎
Proposition 5.6.
The potential wall can not be a totally semistable wall.
Proof.
Thanks to Proposition 5.4, it suffices to show that there is no $w=(r,\Delta,s)$ in $\textbf{{H}}_{W}$ with $\langle w,w\rangle=-2$ satisfying rs<0r-s<0 and $\Re(Z_{\sigma_{t}}(r,\Delta,s)/Z_{\sigma_{t}}(1,0,-1))>0$. Otherwise one has
$r$ $=-\frac{t+9}{40}\Delta\cdot H-\sqrt{\frac{(t+9)^{2}}{400}(\Delta\cdot H)^{2}+4+2\Delta^{2}}$
$s$ $=-\frac{t+9}{40}\Delta\cdot H+\sqrt{\frac{(t+9)^{2}}{400}(\Delta\cdot H)^{2}+4+2\Delta^{2}}$
since $2rs=\Delta^{2}+2$ and $20(r+s)=-(t+9)\Delta\cdot H$.
The condition $\Re(Z_{\sigma_{t}}(r,\Delta,s)/Z_{\sigma_{t}}(1,0,-1))>0$ is equivalent to
$(17t^{2}-3t-4)r+(4t^{2}-2t-2)\Delta\cdot H-5(t+1)s>0$
and, after plug $s$ and $r$, it becomes
$\frac{t-1}{40}(17t^{2}+2t+1)\Delta\cdot H-(17t^{2}+8t+9)\sqrt{\frac{(t+9)^{2}}{400}(\Delta\cdot H)^{2}+4+2\Delta^{2}}>0$
which follows $\Delta\cdot H<0$. Moreover, one has
$\frac{(t-1)^{2}}{1600}(17t^{2}+2t+1)^{2}(\Delta\cdot H)^{2}>(17t^{2}+8t+9)^{2}(\frac{(t+9)^{2}}{400}(\Delta\cdot H)^{2}+4+2\Delta^{2})$
by taking square. The inequality is also
$(\Delta^{2}+2)>\left(\frac{(t+9)^{2}}{800}-\frac{(t-1)^{2}}{3200}\frac{(17t^{2}+2t+1)^{2}}{(17t^{2}+8t+9)^{2}}\right)(\Delta\cdot H)^{2}$
so one has $\Delta^{2}<-2$ and $rs<0$. It means that $r<0$ and $s>0$. Hence
$\frac{(t+9)^{2}}{400}(\Delta\cdot H)^{2}+4+2\Delta^{2}>\frac{(t+9)^{2}}{1600}(\Delta\cdot H)^{2}$
and then
$\frac{3(t+9)^{2}}{3200}(\Delta\cdot H)^{2}>-(2+\Delta^{2})>\left(\frac{(t+9)^{2}}{800}-\frac{(t-1)^{2}}{3200}\frac{(17t^{2}+2t+1)^{2}}{(17t^{2}+8t+9)^{2}}\right)(\Delta\cdot H)^{2}$
which is impossible. ∎
5.3. The wall-crossing of spherical factors
In this subsection, some stable spherical objects will be identified as the Jordan–Hölder factors appearing in the next subsection.
Proposition 5.7.
Consider a line $L\subset S$, then $M_{\sigma_{t}}(0,L,-1)=M_{H}(0,L,-1)$ for any $t\geq 1$.
Proof.
It suffices to show that $M_{\sigma_{t}}(0,L,-1)$ contains only stable objects in the heart of $\sigma_{t}$ for any $t\geq 1$. Otherwise, one has an exact sequence $0\rightarrow A\rightarrow\mathscr{O}_{L}(-2)\rightarrow B\rightarrow 0$ of $\sigma_{t}$-semistable objects with the same phase. Set $v(A)=(r,\Delta,s)$, then $r\geq 1$ and one has
$\Im(Z_{t}(A))=(\frac{1}{5}\Delta\cdot H+\frac{4}{5}r)\sqrt{t}>0\quad\textup{and}\quad\Im(Z_{t}(B))=(\frac{1}{5}-\frac{1}{5}\Delta\cdot H-\frac{4}{5}r)\sqrt{t}>0$
implying that $\Delta\cdot H+4r$ can not be an integer, which is absurd ∎
Proposition 5.8.
Consider a line $L\subset S$, then $M_{\sigma_{t}}(1,-L,0)=M_{H}(1,-L,0)$ for any $t\geq 1$.
Proof.
It suffices to show that $M_{\sigma_{t}}(1,-L,0)$ contains only stable objects in the heart of $\sigma_{t}$ for any $t\geq 1$. Otherwise, one has an exact sequence $0\rightarrow A\rightarrow\mathscr{O}_{S}(-L)\rightarrow B\rightarrow 0$ of $\sigma_{t}$-semistable objects with the same phase such that $A$ is stable. Set $v(A)=(r,\Delta,s)$, then
$\Im(Z_{t}(A))=(\frac{1}{5}\Delta\cdot H+\frac{4}{5}r)\sqrt{t}>0\quad\textup{and}\quad\Im(Z_{t}(B))=(\frac{3}{5}-\frac{1}{5}\Delta\cdot H-\frac{4}{5}r)\sqrt{t}>0$
so that $k:=\Delta\cdot H+4r$ is either $1$ or $2$. It follows $3\Re(Z_{t}(A))=k\Re(Z_{t}(\mathscr{O}_{S}(-L)))$ and then an equation 15s=(3t+12)r4kkt15s=(3t+12)r-4k-kt.
Since (1,L,0)(1,-L,0) is a spherical vector, the stable HN factor $A$ of $\mathscr{O}_{S}(-L)$ is also spherical due to Proposition 3.19 (or directly [11, Lemma 12.2]). So the Hodge index theorem gives
$(4r-k)^{2}\geq 10\Delta^{2}=10(2rs-2)=\frac{4}{3}((3t+12)r-4k-kt)r-20$
and then the inequality $60+3k^{2}\geq t(12r-4kr)r+8kr\geq 12r^{2}+4kr$ as $3r\geq 3>k$. It follows that $r=1$ or $r=2$. The latter case is impossible as one can always see $s\geq 2$ from $t\geq 1$ and the equality for $s$. So the Hodge index theorem becomes $(8-k)^{2}\geq 10\Delta^{2}\geq 60$, impossible.
Hence one must have $r=1$, then $\mathscr{H}^{-1}(B)=0$ from the destabilizing sequence. In this case, the object $B$ is also a spherical stable factor of $\mathscr{O}_{S}(-L)$ by Proposition 3.19. It follows that $s=1-\Delta\cdot L$. One notices that $\Delta\cdot H\leq-2$ and then $s\geq 3$ as $k<3$. But the Hodge index theorem will give $(4-k)^{2}\geq 10\Delta^{2}=10(2s-2)\geq 40$, absurd. ∎
Proposition 5.9.
Consider a conic $C\subset S$, then $M_{\sigma_{t}}(0,C,-1)=M_{H}(0,C,-1)$ for $t\geq 1$.
Proof.
It suffices to show that $M_{\sigma_{t}}(0,C,-1)$ contains only stable objects in the heart of $\sigma_{t}$ for any $t\geq 1$. Otherwise, one has an exact sequence $0\rightarrow A\rightarrow\mathscr{O}_{C}(-2)\rightarrow B\rightarrow 0$ of $\sigma_{t}$-semistable objects with the same phase such that $A$ is $\sigma_{t}$-stable. Set $v(A)=(r,\Delta,s)$, then one has
$\Im(Z_{t}(A))=(\frac{1}{5}\Delta\cdot H+\frac{4}{5}r)\sqrt{t}>0\quad\textup{and}\quad\Im(Z_{t}(B))=(\frac{2}{5}-\frac{1}{5}\Delta\cdot H-\frac{4}{5}r)\sqrt{t}>0$
so that $\Delta\cdot H+4r=1$. Then $2\Re(Z_{t}(A))=\Re(Z_{t}(\mathscr{O}_{C}(-2)))$ and one gets 10s=(2t+8)r510s=(2t+8)r-5.
Since (0,C,1)(0,C,-1) is a spherical vector, the stable factor $A$ is also spherical. So the Hodge index theorem gives $(4r-1)^{2}\geq 10\Delta^{2}=10(2rs-2)=(4t+16)r^{2}-10r-20$ from which one obtains $r\in\{0,1,2\}$ as $r\geq 0$ and $t\geq 1$. Suppose that $r=0$, then one has 2s=12s=-1, which is absurd. Suppose that $r=1$, then $\Delta\cdot H=-3$. However, one can see from $10s=2t+3$ and the Hodge index theorem that $\Delta^{2}=0$, which is impossible as $(S,H)$ is Brill–Noether general of degree $10$. Suppose that $r=2$, then $s\geq 2$ and $\Delta^{2}\geq 6$, violating the Hodge index theorem. ∎
Proposition 5.10.
Consider a conic $C\subset S$, then $M_{\sigma_{t}}(1,-C,0)=M_{H}(1,-C,0)$ for $t\geq 1$.
Proof.
It suffices to show that $M_{\sigma_{t}}(1,-C,0)$ contains only stable objects in the heart of $\sigma_{t}$ for any $t\geq 1$. Otherwise, one has an exact sequence $0\rightarrow A\rightarrow\mathscr{O}_{S}(-C)\rightarrow B\rightarrow 0$ of $\sigma_{t}$-semistable objects with the same phase such that $A$ is $\sigma_{t}$-stable. Set $v(A)=(r,\Delta,s)$, then one has
$\Im(Z_{t}(A))=(\frac{1}{5}\Delta\cdot H+\frac{4}{5}r)\sqrt{t}>0\quad\textup{and}\quad\Im(Z_{t}(B))=(\frac{2}{5}-\frac{1}{5}\Delta\cdot H-\frac{4}{5}r)\sqrt{t}>0$
so that $\Delta\cdot H+4r=1$. Then $\Re(Z_{t}(A))=\Re(Z_{t}(B))$ and one gets 10s=(2t+8)rt410s=(2t+8)r-t-4.
Since (1,C,0)(1,-C,0) is a spherical vector, the stable factor $A$ is also spherical. So the Hodge index theorem gives $(4r-1)^{2}\geq 10\Delta^{2}=10(2rs-2)=(4t+16)r^{2}-(2t+8)r-20$ from which one obtains $r=1$ or $r=2$ as $r\geq 1$ and $t\geq 1$. Suppose that $r=1$, then $\Delta\cdot H=-3$. On the other hand, the Hodge index theorem and $10s=t+4>0$ ensure that $\Delta^{2}=0$. It is impossible since $(S,H)$ is a Brill–Noether general K3 surface. Suppose that $r=2$, then one can see $s\geq 1$ and $\Delta^{2}\geq 6$ but it will violate the Hodge index theorem. ∎
Now one considers the unique Gieseker stable bundle $\mathscr{U}_{S}$ on $S$ with Mukai vector (2,H,3)(2,-H,3) which is known to be the restriction of the tautological bundle $\mathscr{U}_{G}$ on $\operatorname{Gr}(2,V_{5})$.
Proposition 5.11.
One has $M_{\sigma_{t}}(-2,H,-3)=M_{H}(2,-H,3)[1]$ for any $t\geq 1$.
Proof.
It suffices to show that $M_{\sigma_{t}}(-2,H,-3)$ contains only stable objects in the heart of $\sigma_{t}$ for any $t\geq 1$. Otherwise, one has an exact sequence $0\rightarrow A\rightarrow\mathscr{U}_{S}[1]\rightarrow B\rightarrow 0$ of $\sigma_{t}$-semistable objects with the same phase such that $B$ is $\sigma_{t}$-stable. Set $v(B)=(r,\Delta,s)$, then one sees from the destabilizing exact sequence that $r\leq-1$ as $\mathscr{H}^{-1}(B)$ is torsion-free. Moreover, one can see as before that $\Delta\cdot H+4r=1$. It follows $\Re(Z_{t}(B))=\Re(Z_{t}(A))$ and 10s=(2t+8)r+2t710s=(2t+8)r+2t-7.
In this case, both $A$ and $B$ are spherical stable objects as $\mathscr{U}_{S}[1]$ is spherical. So one can get 2s=r22s=r-2. Then r=1r=-1 by the previous equation, which is impossible. ∎
Then one considers the unique Gieseker stable bundle $\mathscr{V}_{S}$ on $S$ with Mukai vector (3,H,2)(3,-H,2) which is actually the restriction of the tautological bundle $\mathscr{V}_{G}$ on $\operatorname{Gr}(3,V_{5})\cong\operatorname{Gr}(2,V_{5})$.
Proposition 5.12.
One has $M_{\sigma_{t}}(3,-H,2)=M_{H}(3,-H,2)$ for any $t\geq 1$.
Proof.
It suffices to show that $M_{\sigma_{t}}(3,-H,2)$ contains only stable objects in the heart of $\sigma_{t}$ for any $t\geq 1$. Otherwise, one has an exact sequence $0\rightarrow A\rightarrow\mathscr{V}_{S}\rightarrow B\rightarrow 0$ of $\sigma_{t}$-semistable objects with the same phase such that $A$ is $\sigma_{t}$-stable. Set $v(A)=(r,\Delta,s)$, then one can see $\Delta\cdot H+4r=1$. It follows $\Re(Z_{t}(A))=\Re(Z_{t}(B))$ and then 10s=(2t+8)r23t10s=(2t+8)r-2-3t.
In this case, both $A$ and $B$ are spherical stable objects as $\mathscr{V}_{S}$ is spherical. So one can get $3s=2r$. Then $s=1$ by the previous equation, which is impossible. ∎
5.4. The flopping walls and identification
In this subsection, we will recover the birational map by identifying the possible flopping walls for the moduli space $\textsf{{M}}_{\sigma_{t}}(v_{1})$.
Proposition 5.13.
Suppose that $t>1$, then the stability condition $\sigma_{t}$ is on a flopping wall for $v_{1}=(1,0,-1)$ if and only if $t=11$ and $S$ contains lines. Moreover, the induced flop map
$S^{[2]}\dashrightarrow Y_{S}$
coincides with the Mukai flop $g$ in Theorem 5.1.
Proof.
According to Theorem 3.18, a flopping wall for it can only be induced by a pairs of vectors $w$ and $w^{\prime}$ in $\textbf{{H}}_{W}$ such that $\langle w,w\rangle=\langle w^{\prime},w^{\prime}\rangle=-2,\langle v_{1},w\rangle=\langle v_{1},w^{\prime}\rangle=1$ and $w^{\prime}+w=v_{1}$. One writes $w=(r,\Delta,s)$, then $(t+9)\Delta\cdot H=-40r+20$ and $\Delta^{2}=2r^{2}-2r-2$. It follows
$20r^{2}-20r-20=10\Delta^{2}\leq(\Delta\cdot H)^{2}=\left(\frac{40r-20}{t+9}\right)^{2}$
or equivalently $r^{2}-r-1\leq 100/((t+9)^{2}-80)$. So $r\in\{1,0\}$ and $t=11$ as $t>1$ and all possible vectors are $(1,\Delta,0)$ with $\Delta^{2}=-2,\Delta\cdot H=-1$ and $(0,\Delta,-1)$ with $\Delta^{2}=-2,\Delta\cdot H=1$.
The divisor class $\Delta$ in $(0,\Delta,-1)$ is in fact the class of a line in $S$ (see, for example, the discussions in [24, Section 2.1.4]) and the class $\Delta$ in $(1,\Delta,0)$ is L-L for some line $L$ in $S$.
According to Theorem 3.14, one has $\textsf{{M}}_{\sigma_{t}}(1,0,-1)=\textsf{{M}}_{H}(1,0,-1)\cong S^{[2]}$ for $t>11$. Then the flopping walls for $\textsf{{M}}_{\sigma_{11}}(1,0,-1)$ are induced by the short exact sequences
$0\rightarrow\mathscr{O}_{S}(-L)\rightarrow\mathscr{I}_{Z}\rightarrow\mathscr{O}_{L}(-Z)\rightarrow 0$
where $Z$ is any length $2$ subscheme on any line $L\subset S$. Then one has a flop
$g\colon S^{[2]}\dashrightarrow Y_{S}=\textsf{{M}}_{\sigma_{t}}(1,0,-1)$
for any $11>t>1$, which coincides with the Mukai flop in Theorem 5.1 by construction. ∎
It remains to show that $Y_{S}\cong X_{S}$ contracts to the moduli space $\textsf{{M}}_{\sigma_{S}}(1,0,-1)$ through the same symplectic resolution $f^{+}\colon X_{S}\rightarrow\tilde{Y}_{A(S)^{\perp}}^{\geq 1}$ as in Theorem 5.1.
Proposition 5.14.
The moduli space $\textsf{{M}}_{\sigma_{S}}(v_{1})$ is isomorphic to $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}$.
Proof.
According to Theorem 3.18, a flopping wall for $\textsf{{M}}_{\sigma_{S}}(v_{1})$ is induced by a pairs of vectors $w$ and $w^{\prime}$ in $\textbf{{H}}_{W}$ such that $\langle w,w\rangle=\langle w^{\prime},w^{\prime}\rangle=-2,\langle v_{1},w\rangle=\langle v_{1},w^{\prime}\rangle=1$ and $w^{\prime}+w=v_{1}$. One writes $w=(r,\Delta,s)$, then $\Delta\cdot H=-4r+2$ and also $\Delta^{2}=2r^{2}-2r-2$. It follows
$20r^{2}-20r-20=10\Delta^{2}\leq(\Delta\cdot H)^{2}=(4r-2)^{2}$
or equivalently $r^{2}-r-1\leq 5$. So one has $r\in\{-2,-1,0,1,2,3\}$.
At first, one notices that $r$ can not be 1-1 or $2$ as in this case one will get a class $\Delta$ with $\Delta\cdot H=6$ and $\Delta^{2}=2$. It is impossible as otherwise $H-\Delta$ would violate Proposition 2.4.
Also, one can see that the class $\Delta$ in $w=(1,-\Delta,0)$ and $w^{\prime}=(0,\Delta,-1)$ with $\Delta\cdot H=2$ and $\Delta^{2}=-2$ is in fact the class of a conic $C$ in $S$ (see for example [24, Section 2.1.4]). The corresponding flopping wall is induced by the short exact sequences
$0\rightarrow\mathscr{O}_{S}(-C)\rightarrow\mathscr{I}_{Z}\rightarrow\mathscr{O}_{C}(-Z)\rightarrow 0$
for the conic $C\subset S$ and any length two subscheme $Z\subset C\subset S$ which is not in the exceptional locus of the Mukai flop $g\colon S^{[2]}\dashrightarrow X_{S}$. So the space $\Sigma_{C}^{\vee}$ is contracted to a point.
Then one consider the pair of vectors $w=(-2,\Delta,-3)$ and $w^{\prime}=(3,-\Delta,2)$ with $\Delta\cdot H=10$ and $\Delta^{2}=10$. The Hodge index theorem ensures that $\Delta=H$. So according to Proposition 5.11 and Proposition 5.12, the corresponding destabilizing exact sequence for $\mathscr{I}_{Z}$ is
$0\rightarrow\mathscr{V}_{S}\rightarrow\mathscr{I}_{Z}\rightarrow\mathscr{U}_{S}[1]\rightarrow 0$
where $Z\subset S$ is a length $2$ subscheme spanning a line not in $S$. Indeed, such a short exact sequence follows from the fact that the ideal sheaf $\mathscr{I}_{L/M_{S}}$ of every line $L\subset M_{S}$ can be written as the cokernel of a unique map $\mathscr{U}_{G}|_{M_{S}}\rightarrow\mathscr{V}_{G}|_{M_{S}}$, and each such map has the ideal sheaf of a line as the cokernel. One just takes the restriction onto $S\subset M_{S}$. The line spanned by $Z$ is not in $S$ because such a length two subscheme $Z$ has been already flopped in Proposition 5.13. So one contracts the Fano variety $F(M_{S})$ of lines, and therefore recovers $f^{+}\colon X_{S}\rightarrow\tilde{Y}_{A(S)^{\perp}}^{\geq 1}$. ∎
Remark 5.15.
Moreover, the involution on $\textsf{{M}}_{\sigma_{S}}(v_{1})$ induced by the group action $\mathfrak{S}^{\vee}_{2}$ coincides with the covering involution on the double dual EPW sextics as the Hodge isometries of the corresponding birational involutions on $X_{S}$ are the same (see [33, Section 6.5] for the generic case). A point-wise description of the involution can be found in [18, Remark 7.8]. Also, the fixed locus $\textsf{{M}}_{\sigma_{S}}(v_{1})_{\mathfrak{S}^{\vee}_{2}}$ of the involution is exactly the ramification locus $Y^{\geq 2}_{A(S)^{\perp}}$.
6. Other EPW varieties as Bridgeland moduli spaces
6.1. Double EPW sextics as moduli spaces
It is shown in [33, Section 6.6] that the moduli space $\textsf{{M}}_{\sigma_{S}}(v_{2})$ is isomorphic to the double EPW sextics once $S$ does not contain any lines. Here one can generalize it using Theorem 1.1 and knowledge on the moduli space and the period of double EPW sextics. To continue, let us recall some notations from [40].
The moduli space of double EPW sextics is defined as the GIT quotient
$\mathfrak{M}_{EPW}:=\operatorname{LG}(\mathsf{\Lambda}^{3}V_{6})/\!/\operatorname{PGL}(V_{6})$
where $\operatorname{LG}(\mathsf{\Lambda}^{3}V_{6})\subset\operatorname{Gr}(10,\land^{3}V)$ is the symplectic Grassmannian parameterizing subspaces which are Lagrangian for the symplectic form given by wedge-product. The period domain $\mathbb{D}_{\Lambda}^{\textup{BB}}$ is defined by the Baily–Borel compactification of the quotient $\mathbb{D}_{\Lambda}$ of a bounded symmetric domain of Type IV by the action of an arithmetic group (see [40, Section 4.2] for details).
Then there exists a rational map
$\mathfrak{p}\colon\mathfrak{M}\dashrightarrow\mathbb{D}_{\Lambda}^{\textup{BB}}$
defined by the hyperkähler period on the open locus where double EPW sextic is a hyperkähler fourfold. According to [40], the rational map is a regular at the open locus $\mathfrak{M}$-$\mathfrak{R}$, where $\mathfrak{R}$ is determined by the Lagrangian subspaces containing decomposable elements.
The mapping $A\mapsto A^{\perp}$ provide an involution $\tau$ on the moduli space $\mathfrak{M}$-$\mathfrak{R}$ which is compatible with the period map $\mathfrak{p}$ i.e. $\mathfrak{p}\circ\tau=\bar{\tau}\circ\mathfrak{p}$ for some involution $\bar{\tau}$ on $\mathbb{D}_{\Lambda}^{\textup{BB}}$.
Proposition 6.1 ([40, Section 4.3 and Theorem 5.1]).
The ramification locus of the involution $\bar{\tau}$ does not intersect with the image $\mathfrak{p}(\mathfrak{M}-\mathfrak{R})$ in $\mathbb{D}_{\Lambda}^{\textup{BB}}$.
Corollary 6.2.
The moduli space $\textsf{{M}}_{\sigma_{S}}(v_{2})$ is isomorphic to the double EPW sextics.
Proof.
The generic case has been verified in [33, Section 6.6], so it remains to prove the statement when $Y^{\geq 3}_{A(S)}\neq 0$. In this case, the special Gushel–Mukai threefold $X$ over $S$ admits a period dual $X^{\prime}$ which is also a special Gushel–Mukai threefold by [31, Lemma 3.8]. Then it follows from [32] that there exists an equivalence ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$. This equivalence descends to an equivalence ${\textbf{{D}}}^{b}(S)\cong{\textbf{{D}}}^{b}(S^{\prime})$ for the strongly smooth Gushel–Mukai surface $S$ associated with the threefold $X^{\prime}$ which is equivariant with the $\mathfrak{S}_{2}^{\vee}$-group actions on both side. So the stability condition $\sigma_{S}$ is sent to $\sigma_{S^{\prime}}$ and the invariant lattice $\mathbb{Z}[v_{1}]\oplus\mathbb{Z}[v_{2}]$ is sent to $\mathbb{Z}[v^{\prime}_{1}]\oplus\mathbb{Z}[v^{\prime}_{2}]$.
In particular, the equivalence either sends $v_{1}$ to $v^{\prime}_{1}$ and $v_{2}$ to $v^{\prime}_{2}$ up to a sign or exchanges them. The first case will lead to $\tilde{Y}^{\geq 1}_{A(S)^{\perp}}\cong\textsf{{M}}_{\sigma_{S}}(v_{1})\cong\textsf{{M}}_{\sigma_{S^{\prime}}}(v^{\prime}_{1})\cong\tilde{Y}^{\geq 1}_{A(S)}$ and then $A(S)\cong A(S)^{\perp}$ due to Theorem 2.10 which will violate the previous proposition. So one concludes. ∎
Remark 6.3.
The parallel results for Gushel–Mukai fourfolds have been established in [21, 42] but only for the very general members.
As a result, one can easily conclude the following corollary using Proposition 4.8 and the identifications of $\textsf{{M}}_{\sigma_{S}}(v_{1})$ and $\textsf{{M}}_{\sigma_{S}}(v_{2})$.
Corollary 6.4.
Consider two special Gushel–Mukai threefolds $X$ and $X^{\prime}$. Then $X$ and $X^{\prime}$ are period partner if and only if one has an equivalence ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$ sending $\kappa_{1}$ to $\kappa^{\prime}_{1}$ and $\kappa_{2}$ to $\kappa^{\prime}_{2}$ up to a sign; $X$ and $X^{\prime}$ are period dual if and only if one has an equivalence ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$ sending $\kappa_{1}$ to $\kappa^{\prime}_{2}$ and $\kappa_{1}$ to $\kappa^{\prime}_{2}$ up to a sign.
6.2. Double dual EPW surfaces as moduli spaces
The statements in this subsection are actually just corollaries of [18, Theorem 7.3] and [28, Theorem 7.12]. The first one describes the associated double dual EPW surfaces, and second one describes the moduli spaces.
Here we only elaborate on the case for a special Gushel–Mukai threefold $X$ over $S$ and the starting point is the Hilbert scheme $C(X)$ of conics on $X$. The surface $C(X)$ has two components $C_{\sigma}(X)$ and $C_{0}(X)$ such that $C_{\sigma}(X)\cong F(M_{S})\cong\mathbb{P}^{2}$ parameterizes the preimage of lines on $M_{S}$. The other component $C_{0}(X)$ is isomorphic to the blow up of the double dual EPW surface associated with $X$ at the distinguished point $\tilde{\bm{p}}_{S}$ according to [18, Theorem 7.3]. These components intersect transversely along the exceptional curve on $C_{0}(X)$.
Due to [28, Theorem 7.12], the moduli space $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})$ is a contraction from $C(X)$ such that the component $C_{\sigma}(X)$ is contracted to the distinguished singular point $\tilde{\bm{p}}_{S}$. It follows that the space $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})$ is nothing but the double dual EPW surface associated with $X$.
Proposition 6.5.
Consider a strongly smooth Gushel–Mukai surface $S$ and the special Gushel–Mukai threefold $X$ over it, then $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})$ is isomorphic to $\tilde{Y}^{\geq 2}_{A(S)^{\perp}}$ and the branched cover
$\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})\rightarrow\textsf{{M}}_{\sigma_{S}}(v_{1})_{\mathfrak{S}^{\vee}_{2}}$
induced by the inflation functor is the natural double cover over $\textsf{{M}}_{\sigma_{S}}(v_{1})_{\mathfrak{S}^{\vee}_{2}}=Y^{\geq 2}_{A(S)^{\perp}}$.
Proof.
The isomorphism $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})\cong\tilde{Y}^{\geq 2}_{A(S)^{\perp}}$ follows from [18, 28] as above. The identification of these branched double covers is similar to [33, Proposition 5.20]. One can do this because the involution on $C(X)$ induced by the branched cover $X\rightarrow M_{S}$ coincides with the one induced by the natural covering involution on $\tilde{Y}^{\geq 2}_{A(S)^{\perp}}$ according to [18, Section 7.3]. In particular, the branched double cover for the double dual EPW surface $\tilde{Y}^{\geq 2}_{A(S)^{\perp}}$ can be described in terms of conics on an open subset and therefore be identified. ∎
To identify double EPW surfaces in the next subsection, we have to use the Kuznetsov component ${\textbf{{Ku}}}(X)$ of an ordinary Gushel–Mukai threefold $X$. Here one recalls that $\textsf{{N}}({\textbf{{Ku}}}(X))$ is also generated by $\kappa_{1}$ and $\kappa_{2}$ as in the special case and there exists a distinguished stability condition $\sigma_{X}$ on ${\textbf{{Ku}}}(X)$ which is sent to $\sigma_{X^{\prime}}$, up to a ${\tilde{\operatorname{GL^{+}}}}(2,\mathbb{R})$-group action, under any equivalence ${\textbf{{Ku}}}(X)\cong{\textbf{{Ku}}}(X^{\prime})$ for any other Gushel–Mukai threefold $X^{\prime}$.
Proposition 6.6.
One has $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})\cong\tilde{Y}^{\geq 2}_{A(X)^{\perp}}$ for an ordinary Gushel–Mukai threefold $X$.
Proof.
It is again a direct application of [18, Theorem 7.3] and [28, Theorem 7.12]. ∎
However, in this case, one does not have enough information to give a double cover as the corresponding moduli space on the associated Gushel–Mukai fourfold is not clear.
6.3. Double EPW surfaces as moduli spaces
Let us start with the easy case.
Corollary 6.7.
Consider a strongly smooth Gushel–Mukai surface $S$ and the special Gushel–Mukai threefold $X$ over it. Suppose in addition that $X$ has a special Gushel–Mukai threefold as period dual, then $\textsf{{M}}_{\sigma_{X}}(-\kappa_{2})$ is isomorphic to $\tilde{Y}^{\geq 2}_{A(S)}$ and the branched cover
$\textsf{{M}}_{\sigma_{X}}(-\kappa_{2})\rightarrow\textsf{{M}}_{\sigma_{S}}(v_{2})_{\mathfrak{S}^{\vee}_{2}}$
induced by the inflation functor is the natural double cover over $\textsf{{M}}_{\sigma_{S}}(v_{2})_{\mathfrak{S}^{\vee}_{2}}=Y^{\geq 2}_{A(S)}$.
Proof.
It follows from Proposition 6.5 by modifying the argument for Proposition 6.2. ∎
A generic special Gushel–Mukai threefold $X$ does not have a special Gushel–Mukai threefold as its period dual according to [31, Lemma 3.8]. On the other hand, one can instead find an ordinary Gushel–Mukai threefold as its period dual by [31, Lemma 3.8]. So one can obtain
Corollary 6.8.
Consider a special Gushel–Mukai threefold $X$, then $\textsf{{M}}_{\sigma_{X}}(-\kappa_{2})\cong\tilde{Y}^{\geq 2}_{A(S)}$.
Proof.
The generic case follows from Proposition 6.6 by modifying the argument for Proposition 6.2. The other cases are covered in Corollary 6.7. ∎
6.4. Description of the special period partners
The stable object $G_{X}$ representing the distinguished singular point $\tilde{p}_{S}$ in $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})$ has the form
$\mathscr{U}_{X}[2]\rightarrow G_{X}\rightarrow\mathscr{V}_{X}^{\vee}[1]\rightarrow\mathscr{U}_{X}[3]$
where $\mathscr{U}_{X}$ and $\mathscr{V}_{X}$ are restrictions of tautological bundles on $\operatorname{Gr}(2,5)$ and $\operatorname{Gr}(3,5)$.
Theorem 6.9 ([29, Theorem 1.1]).
Consider a special Gushel–Mukai threefold $X$ and the special stable object $G_{X}$ in $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})$, then the Brill–Noether locus
$\operatorname{BN}_{X}(G_{X})=\{A\in\textsf{{M}}_{\sigma_{X}}(\kappa_{1}+2\kappa_{2})\,|\,\operatorname{Hom}(A,G_{X}[2])=\mathbb{C}^{\oplus 3}\}\subset\textsf{{M}}_{\sigma_{X}}(\kappa_{1}+2\kappa_{2})$
consists of the object $K_{x}={\textup{{{P}}}}\mathscr{O}_{x}$ for any $x\in X$, where P is the left adjoint functor of the inclusion ${\textbf{{Ku}}}(X)\subset{\textbf{{D}}}^{b}(X)$. Moreover, one has $\operatorname{BN}_{X}(G_{X})\cong X$.
Proof.
One notices that the Kuznetsov component used in [29] is different from ${\textbf{{Ku}}}(X)$ but equivalent. The relation between the distinguished stable object $G_{X}$ and the characterizing object in [29, Theorem 1.1] is explained in [28, Section 5]. The statement about the stable objects in the Brill–Noether locus follows from the argument for [29, Theorem 1.1]. ∎
The representatives for other singular points in the moduli space $\textsf{{M}}_{\sigma_{X}}(-\kappa_{1})$ have been determined in [33, Corollary 6.11]. Each of them is given by a conic in the ramification locus of the covering involution on $X$ and has the form $\mathscr{I}_{C}[1]$. In particular, one has
Corollary 6.10.
A non-trivial special Gushel–Mukai threefold period partner of $X$ corresponding to a conic $C$ is isomorphic to the Brill–Noether locus
$\operatorname{BN}_{X}(\mathscr{I}_{C}[1])=\{A\in\textsf{{M}}_{\sigma_{X}}(\kappa_{1}+2\kappa_{2})\,|\,\operatorname{Hom}(A,\mathscr{I}_{C}[3])=\mathbb{C}^{\oplus 3}\}$
in the moduli space $\textsf{{M}}_{\sigma_{X}}(\kappa_{1}+2\kappa_{2})$. This period partner is denoted by $X_{C}$.
The object $K_{x}$ is in the locus $\operatorname{BN}_{X}(\mathscr{I}_{C}[1])$ if and only if $x\notin C$. So one has a birational map $X_{C}\dashrightarrow X$ whose exceptional locus on $X$ is the conic $C$. It seems that $X_{C}$ is exactly the elementary transformation of $X$ along the conic $C$ (see [27, Section 4.4]).
Acknowledgments
The first author would like to thank his advisors Laura Pertusi and Paolo Stellari for their constant supports, and thank Arend Bayer and Jieao Song for helpful conversations. The authors would like to thank Kieran O’Grady and Sasha Kuznetsov for answering their questions.